### v2

# Exceptional set estimate in prime fields: dimension two implies higher dimensions

Abstract
In this paper, we study in prime fields the exceptional set estimates, which can be viewed as a refinement of Marstrand’s orthogonal projection theorem. Additionally, we address a Furstenberg-type problem, which is closely related. It is shown that the two-dimensional result implies all higher-dimensional results in the prime-field setting.
[…]
[classification=text]
1 Introduction
We study the exceptional set estimate and a Furstenberg-type problem in prime fields. The main novelty of this paper is that we prove, for these two problems, that the two-dimensional result implies all higher-dimensional results.
1.1 Background
We first introduce the two problems—the Furstenberg set problem and the exceptional set estimate in Euclidean space.
1.1.1 Furstenberg set problem
The Furstenberg set problem was originally considered by Wolff [20] in $\mathbb{R}^{2}$. Fix $s\in(0,1],t\in(0,2]$. We say $E\subset\mathbb{R}^{2}$ is an $(s,t)$-set in $\mathbb{R}^{2}$, if $E$ is of the form
$E=\bigcup_{\ell\in\mathcal{L}}Y(\ell),$
where $\mathcal{L}\subset A(1,\mathbb{R}^{2})$ is a set of lines in $\mathbb{R}^{2}$ with $\dim\mathcal{L}=t$ and for each $\ell\in\mathcal{L}$ there exist $Y(\ell)\subset\ell$ with $\dim Y(\ell)=s$. ($\dim$ will always mean Hausdorff dimension.) The Furstenberg set problem asks about the optimal lower bound on the Hausdorff dimension of $(s,t)$-sets, that is,
$\inf_{E\textup{~is~a~}(s,t)\textup{-set}}\dim(E).$
Recently, Ren and Wang [19] resolved the Furstenberg set problem by proving
$\inf_{E\textup{~is~a~}(s,t)\textup{-set}}\dim(E)=\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}.$ (1)
For more references and history on this problem, we refer to [19], [17].
The Furstenberg set problem was also considered in higher dimensions; see, for example, [9], [1], [11]. It can be formulated similarly as follows.
Fix integers $1\leq k<n$. Fix $s\in[0,k],t\in[0,(k+1)(n-k)]$. We say $E\subset\mathbb{R}^{n}$ is a $(s,t;n,k)$-set in $\mathbb{R}^{n}$, if $E$ is of the form
$E=\bigcup_{V\in\mathcal{V}}Y(V),$
where $\mathcal{V}\subset A(k,\mathbb{R}^{n})$ is a set of $k$-planes $\mathbb{R}^{n}$ with $\dim\mathcal{V}=t$ and for each $V\in\mathcal{V}$ there exists $Y(V)\subset V$ with $\dim Y(V)=s$. The Furstenberg set problem in $\mathbb{R}^{n}$ for $k$-planes asks us to compute
$\inf_{E\textup{~is~a~}(s,t;n,k)\textup{-set}}\dim(E).$
The Furstenberg set problem in higher dimensions remains wide open.
Remark 1.1.
The original Furstenberg set problem requires the $k$-planes in $\mathcal{V}$ to have distinct directions. We will not address this version, as it is more challenging than Kakeya conjecture which is open in higher dimensions. Though we still use the term “Furstenberg set problem” to refer to this non-direction-distinct version addressed in this paper, and we hope this does not cause confusion. For progress on the direction-distinct Furstenberg set problem in finite fields, see [2], [3], [4], [5].
1.1.2 Exceptional set estimate
Next, we introduce the problem of the exceptional set estimate. We first introduce this problem over $\mathbb{R}$. The projection theory dates back to Marstrand [12], who showed that if $A$ is a Borel set in $\mathbb{R}^{2}$, then the projection of $A$ onto almost every one-dimensional subspace has Hausdorff dimension $\min\{1,\dim A\}$. This was generalized to $\mathbb{R}^{n}$ by Mattila [13] who showed that if $A$ is a Borel set in $\mathbb{R}^{n}$, then the projection of $A$ onto almost every $k$-dimensional subspace has Hausdorff dimension $\min\{k,\dim A\}$.
One can consider a refined version of Marstrand’s projection problem, which is known as the exceptional set estimate.
For $V\in G(k,\mathbb{R}^{n})$, we use $\pi_{V}:\mathbb{R}^{n}\rightarrow V$ to denote the orthogonal projection. For $A\subset\mathbb{R}^{n}$ and $s>0$, we define the exceptional set
$E_{s}(A;n,k):=\{V\in G(k,\mathbb{R}^{n}):\dim(\pi_{V}(A))<s\}.$
The exceptional set estimate is about finding the optimal upper bound of $\dim(E_{s}(A;n,k))$ (in terms of $\dim A,s,n,k$), that is,
$\sup_{A\subset\mathbb{R}^{n},\dim(A)=a}\dim(E_{s}(A;n,k)).$
For technical reasons, $A$ is always assumed to be Borel.
Much progress has been made on this problem; see, for example, [10], [6], [18], [16], [17].
In $\mathbb{R}^{2}$, D. Oberlin [14] conjectured that for $0<s\leq\min\{1,a\}$ and Borel set $A\subset\mathbb{R}^{2}$ with $\dim A=a$, one has the following exceptional set estimate
$\dim(E_{s}(A;2,1))\leq\max\{2s-a,0\}.$ (2)
This conjecture was proved by Ren and Wang in [19] as a corollary of (1).
1.2 The problems in prime fields
In the remainder of the paper, we look only at the prime fields. Let $p$ be a prime and $\mathbb{F}_{p}$ be the prime field with $p$ elements. Throughout the paper, $A\lessapprox B$ means $A\leq C(\log p)^{C}B$ for some large constant $C$ (independent of $p$). $A\approx B$ means both $A\lessapprox B$ and $B\lessapprox A$. $A\lesssim B$ means $A\leq CB$ for a large constant $C$ (independent of $p$). $A\sim B$ means both $A\lesssim B$ and $B\lesssim A$.
1.2.1 Furstenberg set problem
Definition 1.2 $($s,t;n,k)(s,t;n,k)-set).
Fix integers $1\leq k<n$, and numbers $0\leq s\leq k$, $0\leq t\leq(k+1)(n-k)$. Let $0<\lambda\leq 1$. We say $E$ is a $(s,t;n,k;\lambda)$-Furstenberg set (over $\mathbb{F}_{p}$), if $E$ is a subset of $\mathbb{F}_{p}^{n}$ of form
$E=\bigcup_{V\in\mathcal{V}}Y(V).$
Here, $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{n})$ is a set of $k$-planes with $\mathcal{V}\geq\lambda p^{t}$; for each $V\in\mathcal{V}$, $Y(V)$ is a subset of $V$ with $Y(V)\geq\lambda p^{s}$.
Remark 1.3.
We remark that $s,t,n,k$ are important parameters, while $\lambda$ is not. Since $\lambda$ is not an important parameter, from now on we will drop $\lambda$ from the notation and simply call $(s,t;n,k;\lambda)$-set as $(s,t;n,k)$-set, thinking of $\lambda$ as a fixed constant in $(0,1]$.
We give the following definition.
Definition 1.4.
We say $(s,t;n,k)$ is admissible if $1\leq k<n$ are integers and $0\leq s\leq k$, $0\leq t\leq(k+1)(n-k)$.
The Furstenberg set problem can be formulated as follows:
Question 1.
Suppose $(s,t;n,k)$ is admissible. Find the largest $\mathbf{F}(s,t;n,k)$, so that
$\inf_{E:\ (s,t;n,k)\textup{-set}}\#E\gtrsim_{s,t,n,k,\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon}$ (3)
holds for any $\varepsilon>0$.
In the above inequality, the implicit constant may depend on many parameters, but never depends on $p$. For simplicity, we may view $s,t,n,k$ as fixed parameters and write $\gtrsim_{s,t,n,k,\varepsilon}$ as $\gtrsim_{\varepsilon}$ throughout the paper. The dependence on $s,t,n,k$ will be mentioned when needed. We are interested in the behavior as $p$ goes to infinity, so the goal is to find the index $\mathbf{F}(s,t;n,k)$ for fixed $(s,t;n,k)$.
We see that the problem consists of two parts: for fixed $(s,t;n,k)$,
  1. 1.
    (Lower bound) show for any $(s,t;n,k)$-set $E$, we have $E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon}$;
  2. 2.
    (Upper bound) show the existence of a $(s,t;n,k)$-set $E$, such that $E\lesssim p^{\mathbf{F}(s,t;n,k)}$.
Based on Ren-Wang’s result (1), it is reasonable to make a conjecture in $\mathbb{F}_{p}^{2}$.
Conjecture 1.5.
Fix $s\in(0,1],t\in(0,2]$. Then for any $(s,t;2,1)$-set $E$ over $\mathbb{F}_{p}$, we have
$E\gtrsim_{s,t,\varepsilon}p^{\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}-\varepsilon},$
for any $\varepsilon>0$.
Remark 1.6.
It is crucial to use the prime field $\mathbb{F}_{p}$. Otherwise, the conjecture fails. Consider the problem in $\mathbb{F}_{q}^{2}$, where $q=p^{2}$ is the square of a prime. We choose the set of lines $\mathcal{V}=\{y=ax+b:a,b\in\mathbb{F}_{p}\}$, so $\mathcal{V}=q$ and hence $t=1$. For each line $V:y=ax+b$ in $\mathcal{V}$, define $Y(V):=\{(x,ax+b):x\in\mathbb{F}_{p}\}$, so $Y(V)=p=q^{\frac{1}{2}}$ and therefore $s=\frac{1}{2}$. We can calculate $\big(\bigcup_{V\in\mathcal{V}}Y(V)\big)=p^{2}=q$. However, $\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}=\frac{5}{4}$.
Usually, people believe that the problem in a finite field is easier than in Euclidean space. However, Conjecture 1.5 remains open, whereas the Euclidean version has been resolved.
In the paper, we explicitly find the function $\mathbf{F}(s,t;n,k)$. The definition of $\mathbf{F}(s,t;n,k)$ is tricky, so we postpone it later in Definition 2.1. The following are the main results.
Theorem 1.7 (Upper bound).
Suppose $(s,t;n,k)$ is admissible. Then for any prime $p$, there exits a $(s,t;n,k)$-set $E$ over $\mathbb{F}_{p}$, such that
$E\lesssim p^{\mathbf{F}(s,t;n,k)}.$
Theorem 1.8 (Lower bound).
Suppose $(s,t;n,k)$ is admissible. If we assume Conjecture 1.5 holds, then for any $(s,t;n,k)$-set $E$ over $\mathbb{F}_{p}$, we have
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon},$
for any $\varepsilon>0$.
1.2.2 Exceptional set estimate
We also study the exceptional set estimate in prime fields in all dimensions. We show that the two-dimensional result for the Furstenberg set problem implies all higher-dimensional results for exceptional set estimates.
To define the projection in $\mathbb{F}_{p}^{n}$, we first consider an equivalent definition of the projection in $\mathbb{R}^{n}$.
Recall that $\pi_{V}:\mathbb{R}^{n}\rightarrow V$ is the orthogonal projection. For $V\in G(k,\mathbb{R}^{n})$, we also define the map
$\pi_{V}^{*}:\mathbb{R}^{n}\rightarrow A(k,\mathbb{R}^{n})$
as $\pi^{*}_{V}(x):=x+V$. We see that $\pi_{V}^{*}(x)$ is the unique $k$-plane in $A(k,\mathbb{R}^{n})$ that is parallel to $V$ and passes through $x$. It is not hard to see that $\pi_{V}$ and $\pi^{*}_{V^{\perp}}$ are the same things. This motivates the definition of projections in $\mathbb{F}_{p}^{n}$. We will still use the notations $E_{s}(A;n,k),\pi_{V}^{*}$ for $\mathbb{F}_{p}^{n}$ (which has been used for $\mathbb{R}^{n}$ in the previous subsection). This will not cause confusion.
Definition 1.9 (Exceptional set).
For $V\in G(m,\mathbb{F}_{p}^{n})$, define
$\pi_{V}^{*}:\mathbb{F}_{p}^{n}\rightarrow A(m,\mathbb{F}_{p}^{n})$
as $\pi_{V}^{*}(x):=x+V$.
Suppose $A\subset\mathbb{F}_{p}^{n}$. Let $s>0$. We define
$E_{s}(A;n,k):=\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#\pi_{V}^{*}(A)<p^{s}\}.$
(We remark that $E_{s}(A;n,k)$ also depends on $p$, but for simplicity, we drop $p$ from the notation.)
We explicitly find the function $\mathbf{M}(a,s;n,k)$, whose definition is tricky and therefore postponed to Definition 2.4. We state the following results.
Theorem 1.10 (Lower bound).
Fix $1\leq k<n$, $a\in(0,n]$ and $0<s$. Then for $p$ large enough, there exists $A\subset\mathbb{F}_{p}^{n}$ with $A\gtrsim p^{a}$, such that
$E_{s}(A;n,k)\gtrsim p^{\mathbf{M}(a,s;n,k)}.$ (4)
Theorem 1.11 (Upper bound).
Assume Conjecture 1.5 holds. Fix $1\leq k<n$, $a\in(0,n]$ and $0<s$. If $p$ is large enough, then for any $A\subset\mathbb{F}_{p}^{n}$ with $A\gtrsim p^{a}$, we have
$E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n,k)},$ (5)
for any $\varepsilon>0$.
1.3 Strategy of the proof
The proof of Theorem 1.7 and Theorem 1.10 is by constructing examples. The proof of Theorem 1.8 and Theorem 1.11 is based on a sequence of pigeonhole arguments and induction on both $n$ and $k$. We explain the idea of the proof of Theorem 1.8.
We consider the case $(n,k)=(3,1)$. Suppose that for any $(s^{\prime},t^{\prime};2,1)$-set $E^{\prime}$, we have $E^{\prime}\gtrsim p^{\mathbf{F}(s^{\prime},t^{\prime};2,1)}$. The goal is to prove for any $(s,t;3,1)$-set $E$, $E\gtrsim p^{\mathbf{F}(s,t;3,1)}$. We will carefully analyze the structure of $E$ and find subsets of $E$ which are $(s^{\prime},t^{\prime};2,1)$-set. Then we use the estimate for those $(s^{\prime},t^{\prime};2,1)$-sets.
Write $E=\bigcup_{L\in\mathcal{L}}Y(L)$, where $\mathcal{L}\subset A(1,\mathbb{F}_{p}^{3})$ with $\mathcal{L}=p^{t}$ and $Y(L)\subset L$ with $Y(L)=p^{s}$. Project $\mathcal{L}$ to $\mathbb{F}_{p}^{2}$, and partition $\mathcal{L}$ according to their images. For each $W\in A(1,\mathbb{F}_{p}^{2})$, define $\mathcal{L}_{W}$ to be the set of those $L\in\mathcal{L}$ whose image under the projection $\mathbb{F}_{p}^{3}\rightarrow\mathbb{F}_{p}^{2}$ is $W$. We may assume that all lines in $\mathcal{L}$ are not parallel to $(0,0,1)$ (by throwing away a small portion of $\mathcal{L}$). We obtain the partition
$\mathcal{L}=\bigsqcup_{W\in A(1,\mathbb{F}_{p}^{2})}\mathcal{L}_{W}.$
By pigeonholing, we can find $\mathcal{W}\subset A(1,\mathbb{F}_{p}^{2})$ so that $\mathcal{L}_{W}$ are comparable for all $W\in\mathcal{W}$. Moreover, denoting $\mathcal{W}=p^{t_{1}}$, $\mathcal{L}_{W}=p^{t_{2}}$, we may assume $t_{1}+t_{2}=t$.
Next, we look at
$E_{W}:=\bigcup_{L\in\mathcal{L}_{W}}Y(L).$
This is a $(s,t_{2};2,1)$-set in $W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{2}$. By hypothesis, we have
$E_{W}\gtrsim p^{\mathbf{F}(s,t_{2};2,1)}.$
We write
$E_{W}=\bigsqcup_{\xi\in W}E_{W}\cap(\xi\times\mathbb{F}_{p}).$
By pigeonholing on $E_{W}\cap(\xi\times\mathbb{F}_{p})$, we may assume that there is a subset $\Xi_{W}\subset W$ such that $E_{W}\cap(\xi\times\mathbb{F}_{p})$ are comparable for $\xi\in\Xi_{W}$. Moreover, denoting $\Xi=p^{s_{1}}$, $E_{W}\cap(\xi\times\mathbb{F}_{p})=p^{s_{2}}$, we may assume
$s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};2,1).$
Also, we may assume $s_{1},s_{2}$ are the same for all $W\in\mathcal{W}$ by doing another pigeonholing on $W$. We can also assume a strong condition on $s_{1}$: $s_{1}\geq s$. However, we do not expand the discussion here.
It turns out that the worst scenario for $E_{W}$ is when
$E_{W}=\Xi_{W}\times\{1,\dots,p^{s_{2}}\}.$
This will be made precise in the latter sections, so we do not expand the discussion here.
Now, we have
$E\supset(\bigcup_{W\in\mathcal{W}}\Xi_{W})\times\{1,\dots,p^{s_{2}}\}.$
Note that $\bigcup_{W\in\mathcal{W}}\Xi_{W}$ is a $(s_{1},t_{1};2,1)$-set. Hence,
$E\gtrsim p^{\mathbf{F}(s_{1},t_{1};2,1)+s_{2}}.$
The problem boils down to the following linear programming problem: suppose that
$\begin{cases}t_{1}+t_{2}=t,\\ s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};2,1),\\ s_{1}\geq s.\end{cases}$ (6)
Show that
$\mathbf{F}(s_{1},t_{1};2,1)+s_{2}\geq\mathbf{F}(s,t;3,1).$
1.4 Structure of the paper
In Section 2, we define $\mathbf{F}(s,t;n,k),\mathbf{M}(a,s;n,k)$ and prove some fundamental lemmas. In Section 3, we prove Theorem 1.7. In Sections 4 and 5, we prove Theorem 1.8. In Section 6, we prove Theorem 1.10. In Sections 7 and 8, we prove Theorem 1.11.
2 Preliminary
[…]
Now, we give the definitions of $\mathbf{F}$ and $\mathbf{M}$ that appeared in Theorem 1.7, 1.8, 1.10, 1.11.
Definition 2.1 (Furstenberg index).
Suppose $(s,t;n,k)$ is admissible, we define $\mathbf{F}(s,t;n,k)$ as follows.
(a) If $s=0$, define
$\mathbf{F}(0,t;n,k)=\max\{0,t-k(n-k)\}.$ (7)
(b) If $s>0$, write $s=d+\sigma$ where $d\in\mathbb{N},\sigma\in(0,1]$. If also $t\in[0,(k-d-1)(n-k)]$, define
$\mathbf{F}(s,t;n,k)=s.$ (8)
If we are not in Case (a) and (b), then $s\in(0,k]$ and $t\in((k-d-1)(n-k),(k+1)(n-k)]$, we write
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$ (9)
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $m\in\{0,\dots,n-k-1\}$, $\tau\in(0,d+2]$.
(c) If $\tau\in(0,2]$, define
$\begin{split}\mathbf{F}(s,t;n,k)&=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}\\ &=d+m+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}.\end{split}$ (10)
(d) If $\tau\in(2,d+2]$, define
$\begin{split}\mathbf{F}(s,t;n,k)&=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}\\ &=s+m+1.\end{split}$ (11)
Remark 2.2.
In (c) and (d), $\mathbf{F}$ has the same expression, but we separate them for the convenience of later discussions.
Through this definition, we can calculate $\mathbf{F}(s,t;2,1)=\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}$ for $s\in(0,1],t\in[0,2]$. We see that Conjecture 1.5 is equivalent to saying that $\inf_{E:\ (s,t;2,1)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;2,1)-\varepsilon}$.
The following is a construction of a $(s,t;2,1)$-set, which provides evidence on why Conjecture 1.5 should be reasonable.
Proposition 2.3.
For $0\leq t\leq 2$, there exists a $(0,t;2,1)$-set $E$ such that
$E\lesssim p^{\max\{0,t-1\}}.$ (12)
For $0<s\leq 1,0\leq t\leq 2$.  there exists a $(s,t;2,1)$-set $E$ such that
$E\lesssim p^{\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}}.$ (13)
Proof.
When $s=0$, we construct $E=\bigcup_{L\in\mathcal{L}}Y(L)$ as follows. If $t\leq 1$, choose $\mathcal{L}$ to be a set of $p^{t}$ lines passing through $(0,0)$ and $Y(L)=(0,0)$. Then $E=1=p^{0}$. If $t>1$, we choose $p^{t-1}$ points $E$ on $\mathbb{F}_{p}\times\{0\}$, and for each $x\in E$, we choose p1p-1 lines through $x$ that are not parallel to $\mathbb{F}_{p}\times\{0\}$. If the line $L$ is through $x\in E$, then we let $Y(L)=x$. We see that $E$ is a $(0,t;2,1)$-set with cardinality $p^{t-1}$. (Here, $p^{t-1}$ should actually be the integer $\lceil p^{t-1}\rceil$, but we still write it as $p^{t-1}$ for simplicity, thinking of it as an integer. Same convention applies in the whole paper.)
Consider the case when $s>0$. There are three terms in the “min” on the RHS of (13). We note that: when $t\leq s$, minimum achieves at $s+t$; when $s\leq t\leq 2-s$, minimum achieves at $\frac{3}{2}s+\frac{1}{2}t$; when $2-s\leq t$, minimum achieves at $s+1$.
  1. 1.
    If $E=\bigcup_{L\in\mathcal{L}}Y(L)$ is a $(s,t;2,1)$-set, then $E\leq\sum_{L\in\mathcal{L}}\#Y(L)\leq p^{s+t}$.
  2. 2.
    Choose $\mathcal{L}$ to be a set of $\frac{1}{2}p^{t}$ lines that are not parallel to the line $\mathbb{F}_{p}\times\{0\}$. Let $Y(L)=L\cap(\mathbb{F}_{p}\times\{1,2,\dots,p^{s}\})$. Then $E$ is a $(s,t;2,1)$-set with $E=\bigcup_{L\in\mathcal{L}}Y(L)\subset\mathbb{F}_{p}\times\{1,\dots,p^{s}\}$. Therefore, $E\leq p^{s+1}$.
  3. 3.
    It remains to consider the case when $s+t\leq 2,s\leq t$. This example is of Szemerédi-Trotter type. The heuristic is shown in Figure 1, where we have $\sim p^{\frac{t-s}{2}}$ directions and for each direction, there are $\sim p^{\frac{t+s}{2}}$ lines. We also have a rectangle of size $\sim p^{s}\times p^{\frac{s+t}{2}}$. The rectangle intersects each line $L$ at $\sim p^{s}$ points, which is defined to be $Y(L)$. We also see $E=\bigcup_{L}Y(L)$ is contained in the rectangle. Therefore, $E\lesssim p^{\frac{3}{2}s+\frac{1}{2}t}$.
    Let $\mathcal{L}=\{y=ax+b:a=1,\dots,p^{\frac{t-s}{2}},b=1,\dots,p^{\frac{s+t}{2}}\}$. Here, we view $a,b$ as elements in $\mathbb{F}_{p}$, and hence $y=ax+b$ as a line in $\mathbb{F}_{p}^{2}$. For each $L\in\mathcal{L}$, let $Y(L)=\{(x,y)\in L:x=1,\dots,\frac{1}{2}p^{s}\}$. We see that $E=\bigcup_{L\in\mathcal{L}}Y(L)$ is contained in $\{(x,y):x=1,\dots,\frac{1}{2}p^{s},y=1,\dots,2p^{\frac{s+t}{2}}\}$. Therefore, $E\lesssim p^{\frac{3}{2}s+\frac{1}{2}t}$.
Figure 1:
Next, we define the index for Marstrand’s orthogonal projection.
Definition 2.4 (Marstrand index).
Let $1\leq k<n$ be integers. Let $a\in(0,n],s>0$. We give the following definition for $\mathbf{M}(a,s;n,k)$. We write $a=m+\beta,s=l+\gamma$ where $m,l\in\mathbb{N}$ and $\beta,\gamma\in(0,1]$. For convenience, we call such an expression of $a,s$ canonical.
Type 1: If $s>\min\{a,k\}$, define
$\mathbf{M}(a,s;n,k)=k(n-k).$ (14)
Type 2: If $s\leq\min\{a,k\}$, $l+1\leq m\leq n+l-k$ and $\gamma\in(\beta,1]$, define
$\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}.$ (15)
Type 3: If $s\leq\min\{a,k\}$, $l\leq m\leq n+l-k-1$ and $\gamma\in(0,\beta]$, define
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$ (16)
Type 4: If $s\leq a-(n-k)$, define
$\mathbf{M}(a,s;n,k)=-\infty.$ (17)
Remark 2.5.
$\mathbf{M}(a,s;n,k)=$-$\infty$ corresponds to the case when the exceptional set is empty. We remark that in the Euclidean setting, it is assumed that the empty set has dimension -$\infty$, to satisfy a Fubini-type heuristic: $\dim(A\times B)=\dim A+\dim B$.
One can check, when $(n,k)=(2,1)$,
$\mathbf{M}(a,s;2,1)=\begin{cases}1&s>\min\{1,a\}\\ \max\{0,2s-a\}&a-1<s\leq\min\{1,a\}\\ -\infty&s\leq a-1.\end{cases}$ (18)
We discuss a construction of an exceptional set.
Figure 2:
Proposition 2.6.
Let $a\in(0,2]$, $s\in(\frac{a}{2},\min\{1,a\}]$. Then for $p$ large enough, there exists $A\subset\mathbb{F}_{p}^{2}$ with $A\gtrsim p^{a}$, such that
$E_{s}(A;2,1)\gtrsim p^{2s-a}.$
Proof.
The heuristic is shown in Figure 2. Our set $A$ is the lattice points in the rectangle of size $\frac{1}{5}p^{a-s}\times\frac{1}{5}p^{s}$. For any slope $k\in[1,p^{2s-a}]$, we can find $p^{s}$ lines with slope $k$ to cover $A$. Hence, this slope-$k$ direction belongs to the exceptional set.
Consider the set $A=\{(x,y):|x|\leq\frac{1}{5}p^{a-s},|y|\leq\frac{1}{5}p^{s}\}\subset\mathbb{F}_{p}^{2}$. For $k,m\in\mathbb{F}_{p}$, we use $\ell_{k,m}$ to denote the line $y=kx+m$. We claim that for each $|k|\leq\frac{1}{5}p^{2s-a}$,
$\pi_{\ell_{k,0}}^{*}(A)\subset\{\ell_{k,m}:|m|<\frac{1}{2}p^{s}\}.$
Note that if $(x,y)\in A$ and $|k|\leq\frac{1}{5}p^{2s-a}$, then $|y-kx|\leq\frac{1}{2}p^{s}$, which means that there exists $|m|<\frac{1}{2}p^{s}$ such that $(x,y)\in\ell_{k,m}$. Hence, $\pi_{\ell_{k,0}}^{*}(A)<p^{s}$. This means $E_{s}(A;2,1)\supset\{\ell_{k,0}:|k|\leq\frac{1}{5}p^{2s-a}\}$.
We see that $A\gtrsim p^{a}$, and $E_{s}(A;2,1)$ has cardinality $\gtrsim p^{2s-a}$. ∎
2.2 Useful lemmas
We prove several useful lemmas, which will be used later in the proof. We suggest that the reader skip this part in the first reading.
Lemma 2.7.
Suppose $G(k,\mathbb{F}_{p}^{n})$ is the set of $k$-dimensional subspaces in $\mathbb{F}_{p}^{n}$, $A(k,\mathbb{F}_{p}^{n})$ is the set of $k$-planes in $\mathbb{F}_{p}^{n}$. We have:
  1. 1.
    $G(k,\mathbb{F}_{p}^{n})\sim p^{k(n-k)}$;
  2. 2.
    $A(k,\mathbb{F}_{p}^{n})\sim p^{(k+1)(n-k)}$;
  3. 3.
    Fix integers $l,k$ with $l+k\geq n$. Let $W_{0}\in A(l,\mathbb{F}_{p}^{n})$. Then $\{V\in A(k,\mathbb{F}_{p}^{n}):V\cap W_{0}\textup{~is~a~}(l+k-n)\textup{-plane}\}\gtrsim p^{(k+1)(n-k)}$.
Proof.
(1) and (2) are fundamental. We prove (3). $V\cap W_{0}$ is a (l+kn)(l+k-n)-plane means $V$ intersects $W_{0}$ transversely. (3) agrees with intuition in $\mathbb{R}^{n}$, since generic $V\in A(k,\mathbb{R}^{n})$ intersect $W_{0}$ transversely. Noting (2), we see that it suffices to prove
$\{V\in A(k,\mathbb{F}_{p}^{n}):V\cap W_{0}\textup{~is~a~}r\textup{-plane}\}\lesssim p^{(k+1)(n-k)-1},$ (19)
for r>l+knr>l+k-n. The proof of (19) is as follows. Choose a $r$-plane $L$ in $W_{0}$. There are $G(r,\mathbb{F}_{p}^{l})\sim p^{(r+1)(l-r)}$ choices for $L$. If $V\cap W_{0}=L$, then $V\supset L$. We have $\leq\#\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset L\}=\#G(k-r,\mathbb{F}_{p}^{n-r})\sim p^{(k-r)(n-k)}$ choices for $V$ such that $V\cap W_{0}=L$. The two estimates give (19). ∎
Lemma 2.8.
Let $0\leq l\leq n$, $1\leq k,m\leq n$ be integers that satisfy $n-k\geq m-l$ and $l\leq k,m$. If $W\in G(m,\mathbb{F}_{p}^{n})$, then
$\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(\pi^{*}_{V}(W))\leq p^{l}\}\sim p^{k(n-k)-(k-l)(m-l)}.$
Proof.
By definition
$\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(\pi_{V}^{*}(W))\leq p^{l}\}$ $=\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(V\cap W)\geq\#W/p^{l}=p^{m-l}\},$
The latter set is
$\bigcup_{j\geq m-l}\bigcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}.$
It suffices to show for $j\geq m-l$,
$\bigcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}\sim p^{k(n-k)-(k+j-m)j}.$ (20)
On the one hand, the LHS is
$\begin{split}&\leq\sum_{U\in G(j,W)}\#\{V\in G(n-k,n):V\supset U\}=\#G(j,\mathbb{F}_{p}^{m})\#G(n-k-j,\mathbb{F}_{p}^{n-j})\\ &\sim p^{j(m-j)}p^{(n-k-j)k}=p^{k(n-k)-(k+j-m)j}.\end{split}$ (21)
On the other hand, for fixed $U\in G(j,W)$, we have
$\begin{split}&\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\\ \supset&\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}\setminus\bigcup_{j+1\leq j^{\prime}\leq n-k}\bigcup_{U^{\prime}\in G(j^{\prime},W),U^{\prime}\supset U}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U^{\prime}\}\end{split}$ (22)
One can check $\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}=\#G(n-k-j,\mathbb{F}_{p}^{n-j})=p^{k(n-k-j)}$. $\bigcup_{j+1\leq j^{\prime}\leq n-k}\bigcup_{U^{\prime}\in G(j^{\prime},W),U^{\prime}\supset U}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U^{\prime}\}$ has cardinality
$\begin{split}&\leq\sum_{j+1\leq j^{\prime}\leq n-k}\#G(j^{\prime}-j,\mathbb{F}_{p}^{m-j})\#G(n-k-j^{\prime},n-j^{\prime})\\ &\sim\sum_{j+1\leq j^{\prime}\leq n-k}p^{(j^{\prime}-j)(m-j^{\prime})+(n-k-j^{\prime})k}\\ &=\sum_{j+1\leq j^{\prime}\leq n-k}p^{k(n-k-j)+(j^{\prime}-j)(m-k-j^{\prime})}\\ &\lesssim p^{k(n-k-j)-1}.\end{split}$ (23)
In the last line, we use $(j^{\prime}-j)(m-k-j^{\prime})\leq m-k-j-1\leq m-k-(m-l)-1\leq-1$. Therefore,
$\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\sim p^{k(n-k-j)}.$
We can bound the left-hand side of (20) by
$\geq\#\bigsqcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\sim p^{j(m-j)+k(n-k-j)}=p^{k(n-k)-(k+j-m)j}.$
Lemma 2.9.
$\mathbf{F}(s,t;n,k)$ is increasing in $s$ and $t$.
Lemma 2.10.
Suppose $(s,t;n,k)$ is admissible, then
$\mathbf{F}(s,t;n,k)\geq s+\max\{0,t-k(n-k)\}.$
Proof.
It is easy to check $\mathbf{F}(s,t;n,k)\geq s$, so we just need to check $\mathbf{F}(s,t;n,k)\geq s+t-k(n-k)$. We check four cases as in Definition 2.1.
(a) If $s=0$, this is true by definition.
(b) If $s>0$, write $s=d+\sigma.$ If $t\in[0,(k-d-1)(n-k)]$, then
$\mathbf{F}(s,t;n,k)=s\geq s+t-k(n-k).$
(c) If (9) holds and $\tau\in(0,2]$, then
$\mathbf{F}(s,t;n,k)=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
To show it is $\geq s+t-k(n-k)$ is equivalent to show
$(d+1)(n-k-m)+\min\{\tau,\frac{1}{2}\sigma+\frac{1}{2}\tau,1\}\geq\tau.$
Since $(d+1)(n-k-m)\geq 1\geq\frac{1}{2}\tau$ and $\min\{\tau,\frac{1}{2}\sigma+\frac{1}{2}\tau,1\}\geq\frac{1}{2}\tau$, it is proved.
(d) In the final case, we have
$\mathbf{F}(s,t;n,k)=s+m+1.$
To show it is $\geq s+t-k(n-k)$ is equivalent to show
$(d+1)(n-k-m)+1\geq\tau.$
This is true since $(d+1)(n-k-m)\geq d+1$ and $\tau\leq d+2$. ∎
Lemma 2.11.
$\mathbf{F}(s,t;n,k)$ is $1$-Lipschitz in $t$. More precisely, suppose that $(s,t_{1}+t_{2};n,k)$ is admissible, $t_{1},t_{2}\geq 0$. Then
$\mathbf{F}(s,t_{1}+t_{2};n,k)\leq t_{1}+\mathbf{F}(s,t_{2};n,k).$
Proof.
This is equivalent to showing that for fixed $s,n,k$, the map $t\mapsto\mathbf{F}(s,t;n,k)$ is Lipschitz with Lipschitz constant $\leq 1$. We sketch the proof. First, it is not hard to show that $t\mapsto\mathbf{F}(s,t;n,k)$ is continuous. Since $\mathbf{F}(s,t;n,k)$ is piecewise defined, we check each formula of $\mathbf{F}(s,t;n,k)$. We see that the derivative on $t$ is $\leq 1$. ∎
$\mathbf{F}(s,t;n,k)$ is continuous in $t$, but not in $s$. However, we can still say something about $s$.
Lemma 2.12.
$\mathbf{F}(s,t;n,k)$ is locally left-Lipschitz in $s$. In other words, if we write $s=d+\sigma$ ($d\in\mathbb{N},\sigma\in(0,1]$), we have
$\mathbf{F}(s-\theta,t;n,k)\geq\mathbf{F}(s,t;n,k)-O(\theta)$
for any $0\leq\theta<\sigma$.
Proof.
We just note that $s\mapsto\mathbf{F}(s,t;n,k)$ is defined piecewise for $s$ in left-open intervals of the form $(d,d+1]$ where $d\in\mathbb{N}$. Moreover, $s\mapsto\mathbf{F}(s,t;n,k)$ is Lipschitz in each of these left-open intervals. ∎
The next results are for the Marstrand index.
Lemma 2.13.
Each pair $(a,s)$ satisfying $a\in(0,n],s>0$ belongs to exactly one of the four types in Definition 2.4.
Proof.
To show this, we just need to prove that for a fixed $a\in(0,n]$, the ranges of $s$ determined by four cases form a partition of $(0,\infty)$. Type 1 corresponds to $s\in(\min\{a,k\},\infty)$. Type 4 corresponds to $s\in(0,a-(n-k)]$. Note that $a-(n-k)\leq\min\{a,k\}$, so we just need to show the ranges of $s$ determined by Type 2 and Type 3 from a partition of $(a-(n-k),\min\{a,k\}]$. Since the ranges of $s$ determined by Type2 and Type 3 are disjoint, we just need to show
$\{s:s\textup{~satisfies~Type~}2\}\bigcup\{s:s\textup{~satisfies~Type~}3\}=(a-(n-k),\min\{a,k\}].$
We first show $\subset$". Suppose $s$ satisfies the condition in Type 2. We want to show s>a(nk)s>a-(n-k), which is equivalent to $m<n+l-k+\gamma-\beta$. This is true since $m\leq n+l-k$ and $\gamma>\beta$. Suppose $s$ satisfies the condition in Type 3. Again, we want to show $m<n+l-k+\gamma-\beta$. This is true since $m\leq n+l-k-1$ and $\gamma-\beta>-\beta\geq-1$.
Next, we show $\supset$”. Suppose $s=l+\gamma\in(a-(n-k),\min\{a,k\}]$. s>a(nk)s>a-(n-k) implies
$m<n+l-k+\gamma-\beta.$
$s\leq a$ implies
$m\geq l+\gamma-\beta.$
If $\gamma\in(\beta,1]$, then $l+1\leq m\leq n+l-k$. Hence, $s$ satisfies the condition in Type 2. If $\gamma\in(0,\beta]$, then $l\leq m\leq n+l-k-1$. Hence, $s$ satisfies the condition in Type 3. ∎
Lemma 2.14.
$\mathbf{M}(a,s;n,k)$ is decreasing in $a$ and increasing in $s$.
Lemma 2.15.
For $(a,s,n,k)$ and $(a-\theta,s-\theta;n,k)$ that are in the domain of $\mathbf{M}$ and $\theta\geq 0$, we have
$\mathbf{M}(a-\theta,s-\theta;n,k)\leq\mathbf{M}(a,s;n,k).$
Proof.
Let $a=m+\beta,s=l+\gamma$ be the canonical expression. We just check four types as in Definition 2.4.
Type 1: If $s>\min\{a,k\}$, then
$\mathbf{M}(a-\theta,s-\theta;n,k)\leq k\leq\mathbf{M}(a,s;n,k).$
Type 4: If $s\leq a-(n-k)$, then we have $\mathbf{M}(a-\theta,s-\theta;n,k)=\mathbf{M}(a,s;n,k)=$-$\infty$.
Type 2: Suppose $(a,s;n,k)$ is of Type 2.
If $\theta<\beta$.
$\mathbf{M}(a-\theta,s-\theta;n,k)=k(n-k)-(m-l)(k-l)+\max\{2(\gamma-\theta)-(\beta-\theta+1),0\}$
$\leq k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}=\mathbf{M}(a,s;n,k).$
If $\beta\leq\theta<\gamma$, then $a-\theta=(m-1)+(1+\beta-\theta),s-\theta=l+(\gamma-\theta)$. So, $(a-\theta,s-\theta;n,k)$ is of Type 3. We have the same estimate:
$\mathbf{M}(a-\theta,s-\theta;n,k)=k(n-k)-(m-l)(k-l)+\max\{2(\gamma-\theta)-(\beta-\theta+1),0\}$
$\leq k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}=\mathbf{M}(a,s;n,k).$
Note that the integer part $(m,l)$ is reduced to (n1,l)(n-1,l), so we can continue doing the iteration.
Type 3: This is similar to Type 2, so we just omit the proof.
Lemma 2.16.
Suppose $a-(n-k)<s\leq\min\{k,a\}$. Write $a=m+\beta,s=l+\gamma$ as in Definition 2.4. Then
$\mathbf{M}(a,s;n,k)\leq k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\},$
and
$\mathbf{M}(a,s;n,k)\geq k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
Proof.
The proof is simply by checking when $(a,s;n,k)$ is of Type 2 or Type 3. ∎
We also show two easy results for the exceptional set estimate.
Lemma 2.17.
Recall Definition 2.4. Suppose that we are in Type 1: $s>\min\{k,a\}$. Then there exists $A\subset\mathbb{F}_{p}^{n}$ with $A\sim p^{a}$, such that
$E_{s}(A;n,k)\gtrsim p^{k(n-k)}.$
Proof.
We just choose $A$ to be any set with cardinality $p^{a}$. One can check by definition that $E_{s}(A;n,k)=G(k,\mathbb{F}_{p}^{n})$. ∎
Lemma 2.18.
Recall Definition 2.4. Suppose that we are in Type 4: $s\leq a-(n-k)$. Then for any $A\subset\mathbb{F}_{p}^{n}$ with $A\geq p^{a}$, we have
$E_{s}(A;n,k)=0.$
Proof.
If $V\in E_{s}(A;n,k)$, then
$p^{a}=\#A=\sum_{L\in\pi_{V}^{*}(A)}\#(L\cap A)\leq\#\pi_{V}^{*}(A)\cdot\#V<p^{s+n-k},$
a contradiction! Therefore, $E_{s}(A;n,k)=\emptyset$. ∎
3 Furstenberg set problem: upper bound
In this section, we prove Theorem 1.7. In other words, we will construct a $(s,t;n,k)$-set $E=\bigcup_{V\in\mathcal{V}}Y(V)$, such that
$E\lesssim_{s,t,n,k}p^{\mathbf{F}(s,t;n,k)}.$
3.1 Two special cases
To familiarize ourselves with $\mathbf{F}(s,t;n,k)$, we discuss two special cases: $s=k$; $k=1$.
When $s=k$, we can morally view $Y(V)$ as $V$. Therefore, a $(k,t;n,k)$-set is of the form
$E=\bigcup_{V\in\mathcal{V}}V.$
Estimating the size of $E$ is known as the problem of “union of $k$ planes”. It has been resolved in finite fields by R. Oberlin [15]. The Euclidean version was resolved by the author in [7].
When $k=1$, this is the Furstenberg set problem for lines. To construct a $(s,t;n,1)$-set
$E=\bigcup_{L\in\mathcal{L}}Y(L)$
with size as small as possible, we hope that the lines in $\mathcal{L}$ are compressed in a low-dimensional space. Write $t=2m+\beta$, where $0\leq m\leq n-2$ is an integer, and $\beta\in(0,2]$. Since $t\leq 2(m+1)$, we can make $\mathcal{L}$ contained in $\mathbb{F}_{p}^{m+2}$. Therefore, our goal becomes to construct a small $(s,t;m+2,1)$-set. The trick is to take the Cartesian product of a $(s,\beta;2,1)$-set with $\mathbb{F}_{p}^{m}$. Let $E^{\prime}\subset\mathbb{F}_{p}^{2}$ be a $(s,\beta;2,1)$-set and let $E^{\prime}$ have the form
$E^{\prime}=\bigcup_{L^{\prime}\in\mathcal{L}^{\prime}}Y(L^{\prime}).$
We will construct a $(s,t;m+2,1)$-set $E$ based on $E^{\prime}$. Let $\pi_{\mathbb{F}_{p}^{2}}:\mathbb{F}_{p}^{m+2}\rightarrow\mathbb{F}_{p}^{2}$ be the projection onto the first two coordinates. Define $\mathcal{L}$ to be the set of lines $L$ in $\mathbb{F}_{p}^{m+2}$ such that $\pi_{\mathbb{F}_{p}^{2}}(L)$ is a line in $\mathcal{L}^{\prime}$. Then, $\mathcal{L}\sim\#G(1,\mathbb{F}_{p}^{m+1})\cdot\#\mathcal{L}^{\prime}\sim p^{2m+\beta}=p^{t}$. If $L\in\mathcal{L}$, then $L^{\prime}=\pi_{\mathbb{F}_{p}^{2}}(L)\in\mathcal{L}^{\prime}$. We define $Y(L):=L\cap\pi_{\mathbb{F}_{p}^{2}}^{-1}(Y(L^{\prime}))$. We see that $Y(L)=\#Y(L^{\prime})\sim p^{s}$, so
$E=\bigcup_{L\in\mathcal{L}}Y(L)$
is a $(s,t;m+2,1)$-set. Also, since $E\subset E^{\prime}\times\mathbb{F}_{p}^{m}$, we have
$E\leq\#E^{\prime}\cdot p^{m}\lesssim p^{m+\min\{s+\beta,\frac{3}{2}s+\frac{1}{2}\beta,s+1\}}$
By checking Definition 2.1, one exactly sees that
$m+\min\{s+\beta,\frac{3}{2}s+\frac{1}{2}\beta,s+1\}=\mathbf{F}(s,t;m+2,1)=\mathbf{F}(s,t;n,1).$
3.2 Proof of Theorem 1.7
For admissible $(s,t;n,k)$, we will construct a $(s,t;n,k)$-set
$E=\bigcup_{V\in\mathcal{V}}Y(V)$
such that
$E\gtrsim p^{\mathbf{F}(s,t;n,k)}.$
[…]
If $t\geq k(n-k)$, we choose $p^{t-k(n-k)}$ points $X$ in $\mathbb{F}_{p}^{n-k}$. For each such point $x\in X$, we choose $\sim p^{k(n-k)}$ $k$-planes that pass through $x$, and require the $k$-planes to be transverse to $\mathbb{F}_{p}^{n-k}$. We obtain $\sim p^{t}$ $k$-planes which is $\mathcal{V}$. If $V$ passes through $x\in X$, we set $Y(V)=\{x\}$. We see that $E=\#X\leq p^{t-k(n-k)}$.
Case (b): $s>0,t\in[0,(k-d-1)(n-k)]$
We write $s=d+\sigma$, where $d\in\{0,\dots,k-1\},\sigma\in(0,1]$. The construction is similar to Case (a). We choose $\sim p^{t}$ many $k$-planes that contain $\mathbb{F}_{p}^{d+1}$. This is doable since
$\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset L\}=\#G(k-d-1,\mathbb{F}_{p}^{n-d-1})\sim p^{(k-d-1)(n-k)}.$
Let $\mathcal{V}$ be these $k$-planes. Fix $Y\subset\mathbb{F}_{p}^{d+1}$ with $Y=p^{s}$. For each $V\in\mathcal{V}$, let $Y(V)=Y$. We see that $E=\#Y=p^{s}$.
Case (c): $s\in(0,k],t\in((k-d-1)(n-k),(k+1)(n-k)]$
We write
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$ (24)
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $m\in\{0,\dots,n-k-1\}$, $\tau\in[0,2]$. We remark that here we allow $\tau=0$, although we assumed $\tau>$ in the canonical expression. Allowing $\tau=0$ will produce the example in Case (d).
We write $\mathbb{F}_{p}^{n}=\mathbb{F}_{p}^{n-k+d+1}\times\mathbb{F}_{p}^{k-d-1}$. Note that $\gtrsim 1$ fraction of $k$-planes intersect $\mathbb{F}_{p}^{n-k+d+1}$ at a $(d+1)$-plane. Heuristically speaking, to make #E\#E as small as possible, we would like $E$ to be a subset of $\mathbb{F}_{p}^{n-k+d+1}\times\mathbf{0}^{k-d-1}$. For simplicity, we write $\mathbb{F}_{p}^{n-k+d+1}\times\mathbf{0}^{k-d-1}$ as $\mathbb{F}_{p}^{n-k+d+1}$, thinking of it as a subspace of $\mathbb{F}_{p}^{n}$. This will not cause any ambiguity.
Note that for each $W\in A(d+1,\mathbb{F}_{p}^{n-k+d+1})$, the cardinality of
$\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset W,V\textup{~is~transverse~to~}\mathbb{F}_{p}^{n-k+d+1}\}$
is $\sim\#G(k-d-1,\mathbb{F}_{p}^{n-k})\sim p^{(k-d-1)(n-k)}$. We just need to find $\mathcal{W}\subset A(d+1,\mathbb{F}_{p}^{n-k+d+1})$ with $\mathcal{W}\sim p^{t-(k-d-1)(n-k)}=p^{(d+2)m+\tau}$, and for each $W\in\mathcal{W}$, to find $Y(W)\subset W$ with $Y(W)\sim p^{s}$, such that
$\bigcup_{W\in\mathcal{W}}Y(W)\lesssim p^{\mathbf{F}(s,t;n,k)}.$ (25)
If this is done, then we let $\mathcal{V}$ be those $k$-planes $V$ that contain some $W\in\mathcal{W}$, and $Y(V)=Y(W)$ if $V\supset W$. We see that $\mathcal{V}\sim\#\{W\}\cdot p^{(k-d-1)(n-k)}\sim p^{t}$, and hence $E=\bigcup_{V\in\mathcal{V}}Y(V)$ is a $(s,t;n,k)$-set. Also, we have the upper bound
$E=\#\bigcup_{W\in\mathcal{W}}Y(W)\lesssim p^{\mathbf{F}(s,t;n,k)}.$
It suffices to find a $(s,(d+2)m+\tau;n-k+d+1,d+1)$-set $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ such that (25) holds. Note that $A(d+1,\mathbb{F}_{p}^{m+d+2})\sim p^{(d+2)(m+1)}$ and $(d+2)m+\tau\leq(d+2)(m+1)$. We will construct the $(d+1)$-planes $\mathcal{W}$ so that they lie in $\mathbb{F}_{p}^{m+d+2}$. It suffices to find a $(s,(d+2)m+\tau;m+d+2,d+1)$-set $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ such that (25) holds.
We write $\mathbb{F}_{p}^{m+d+2}=\mathbb{F}_{p}^{d+2}\times\mathbb{F}_{p}^{m}=\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{d}\times\mathbb{F}_{p}^{m}$. We first choose a $(\sigma,\tau;2,1)$-set $E_{2}=\bigcup_{L\in\mathcal{L}}Y(L)$ in $\mathbb{F}_{p}^{2}$, such that
$E_{2}\lesssim p^{\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
The existence of $E_{2}$ was shown in Proposition 2.3.
Next, we define $\mathcal{U}:=\{L\times\mathbb{F}_{p}^{d}:L\in\mathcal{L}\}$ which is a subset of $A(d+1,\mathbb{F}_{p}^{d+2})$. And for each $U=L\times\mathbb{F}_{p}^{d}\in\mathcal{U}$, define $Y(U)=Y(L)\times\mathbb{F}_{p}^{d}$. We see that $\mathcal{U}=\#\mathcal{L}\sim p^{\tau}$, $Y(U)=\#Y(L)\cdot p^{d}\sim p^{\sigma+d}=p^{s}$. We also see that $E_{3}:=\bigcup_{U\in\mathcal{U}}Y(U)$ is contained in $E_{2}\times\mathbb{F}_{p}^{d}$, and hence
$E_{3}\lesssim p^{d+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
Finally, we define $\mathcal{W}$. For each $U\in\mathcal{U}$, consider the $(m+d+1)$-plane $U\times\mathbb{F}_{p}^{m}$. Let $\pi_{U}:U\times\mathbb{F}_{p}^{m}\rightarrow U$ be the projection (naturally defined). We let $\mathcal{W}_{U}$ be the set of $(d+1)$-planes $W$ contained in $U\times\mathbb{F}_{p}^{m}$ such that $\pi_{U}(W)=U$. Note that $\gtrsim 1$ fraction of the $(d+1)$-planes in $U\times\mathbb{F}_{p}^{m}$ lie in $\mathcal{W}_{U}$, so $\mathcal{W}_{U}\sim\#A(d+1,\mathbb{F}_{p}^{m+d+1})\sim p^{(d+2)m}$.
We define $\mathcal{W}=\bigcup_{U\in\mathcal{U}}\mathcal{W}_{U}$ (actually this is a disjoint union). For $W\in\mathcal{W}_{U}$, define $Y(W)=W\cap\pi_{U}^{-1}(Y(U))$. We see that $\mathcal{W}\sim\#\mathcal{U}\cdot p^{(d+2)m}\sim p^{(d+2)m+\tau}$, and $Y(W)=\#Y(U)\sim p^{s}$. Hence, $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ is a $(s,(d+2)m+\tau;m+d+2,d+1)$-set. Since $E_{1}\subset E_{3}\times\mathbb{F}_{p}^{m}$, we have
$E_{1}\leq\#E_{3}\cdot p^{m}\lesssim p^{d+m+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
The construction is done.
Case (d)
We still use (24), but the range of $\tau$ is $\tau\in(2,d+2]$. Note that the worst case is when $\tau=d+2$. What we need to do is for
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)(m+1),\end{cases}$ (26)
construct a $(s,t;n,k)$-set $E$ such that
$E\lesssim p^{s+m+1}.$
The trick is to still use the construction in Case (c), with $\tau$ replaced by 0 and $m$ replaced by $m+1$. Therefore, we can construct a $(s,t;n,k)$-set $E$ such that
$E\lesssim p^{d+m+1+\min\{\sigma+0,\frac{3}{2}\sigma+\frac{1}{2}0,\sigma+1\}=p^{d+m+1+\sigma}=p^{s+m+1}}.$
4 Furstenberg set problem: Lower bound I
To become more familiar with the problem, we first prove two special cases of Theorem 1.8: $s=0$; $t=0$.
Proposition 4.1.
If $E$ is a $(0,t;n,k)$-set, then
$E\gtrsim p^{\mathbf{F}(0,t;n,k)}=p^{\max\{0,t-k(n-k)\}}.$
Proof.
Suppose $E=\bigcup_{V\in\mathcal{V}}Y(V)$. Note that each $x\in E$ can lie in at most $G(k,\mathbb{F}_{p}^{n})\sim p^{k(n-k)}$ different $V$. We have
$E\cdot p^{k(n-k)}\geq\sum_{V\in\mathcal{V}}\#Y(V)\gtrsim p^{t}.$
Hence, $E\gtrsim p^{t-k(n-k)}$. Also, trivially, $E\geq 1$. ∎
Proposition 4.2.
If $E$ is a $(s,0;n,k)$-set, then
$E\gtrsim p^{\mathbf{F}(s,0;n,k)}=p^{s}.$
Proof.
By the definition of $(s,0;n,k)$-set, $E\geq\#Y(V)\gtrsim p^{s}$. ∎
In this section, we prove Theorem 1.8 for $(n,k)=(k+1,k)$. The proof is by induction on $k$. The goal of this section is to prove the following result.
Proposition 4.3.
Let $k\in\mathbb{N}$ and $k\geq 2$. Suppose for admissible (s,t;k,k1)(s,t;k,k-1), we know
$\inf_{E:\ (s,t;k,k-1)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k,k-1)-\varepsilon}$ (27)
for any $\varepsilon>0$. Then for admissible $(s,t;k+1,k)$, we have
$\inf_{E:\ (s,t;k+1,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k+1,k)-\varepsilon},$ (28)
for any $\varepsilon>0$.
We mention a technical thing in the proof. For $\eta>0$, denote $\eta\mathbb{Z}:=\{\eta n:n\in\mathbb{Z}\}$. In Theorem 1.8, the implicit constant in the inequality also depends on $s,t$, and therefore may not be uniformly bounded. To get rid of it, we choose a small $\eta>0$, and discretize $s,t$, so that $s,t\in\eta\mathbb{Z}$. Then we obtain the bound
$E\gtrsim_{\varepsilon,\eta,n,k}p^{\mathbf{F}(s-\eta,t-\eta;n,k)-\varepsilon},$
where the implicit constant is independent of $s,t$.
Lemma 4.4.
Assume (27) holds. Suppose $(s,t;k+1,k)$ is admissible. Fix $0<\varepsilon<1/100$. Then for any $(s,t;k+1,k)$-set $E$, there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
$\begin{cases}1\leq p^{t_{1}}\leq p^{k-1},1\leq p^{t_{2}}\leq p^{2};\\ 1\leq p^{s_{1}}\leq p,1\leq p^{s_{2}}\leq p^{s};\\ p^{s_{1}}\lesssim p^{u}\leq p,1\leq p^{v}\leq p;\end{cases}$ (29)
and
$\begin{cases}p^{t_{1}+t_{2}}\gtrapprox p^{t};\\ p^{s_{1}+s_{2}}\gtrapprox p^{s};\\ p^{u+v}\gtrsim_{\varepsilon}p^{\mathbf{F}((s_{1}-\varepsilon^{2})_{+},(t_{2}-\varepsilon^{2})_{+};2,1)-\varepsilon};\end{cases}$ (30)
and
$E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}((s_{2}-\varepsilon^{2})_{+},(t_{1}+v-\varepsilon^{2})_{+};k,k-1),s_{2}+v\}-\varepsilon}.$ (31)
Here, $(x)_{+}=\max\{0,x\}$. In addition, the implicit constants do not depend on $t_{1},t_{2},s_{1},s_{2},u,v$.
Proof.
We will do several steps of pigeonholings, which lead to the parameters $t_{1},t_{2},s_{1},s_{2},u,v$.
By the definition of $(s,t;k+1,k)$-set, we write
$E=\bigcup_{V\in\mathcal{V}}Y(V),$
where $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{k+1})$ with $\mathcal{V}\sim p^{t}$, and $Y(V)\subset V$ with $Y(V)\sim p^{s}$.
The strategy is to find a subset of $\mathcal{V}$, and a subset of $Y(V)$ for each $V$, still denoted by $\mathcal{V}$, $Y(V)$, so that the new $\mathcal{V},Y(V)$ satisfy certain uniformity. This is done by a sequence of pigeonhole arguments.
By properly choosing the coordinate, we can assume $\gtrsim 1$ fraction of the $k$-planes in $\mathcal{V}$ are not parallel to $\mathbb{F}_{p}^{k}$. Throwing away some $V\in\mathcal{V}$, we can just assume all the $k$-planes in $\mathcal{V}$ are not parallel to $\mathbb{F}_{p}^{k}$.
Now, we partition $\mathcal{V}$ according to $G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$. Since each $V\in\mathcal{V}$ is not parallel to $\mathbb{F}_{p}^{k}$ and the ambient space is $\mathbb{F}_{p}^{k+1}$, $V$ must intersect $\mathbb{F}_{p}^{k}\times\mathbf{0}$ at a (k1)(k-1)-plane. For each $W\in G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$, we define $\mathcal{V}_{W}$ to be the set of $V\in\mathcal{V}$ such that $V\cap\mathbb{F}_{p}^{k}$ is parallel to $W$. We obtain a partition
$\mathcal{V}=\bigsqcup_{W\in G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})}\mathcal{V}_{W}.$
First pigeonholing. We assume that there exists $\mathcal{W}\subset G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$ with $\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, $\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $1\leq p^{t_{1}}\leq p^{k-1},1\leq p^{t_{2}}\leq p^{2}$ are dyadic numbers, and
$p^{t_{1}}p^{t_{2}}\gtrapprox p^{t}.$ (32)
(Recall that $A\gtrapprox B$ means $A\gtrsim(\log p)^{C}B$.) We still use $\mathcal{V}$ to denote $\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$. And we see that
$\mathcal{V}\gtrapprox p^{t}.$
Second pigeonholing. By properly choosing the coordinate and throwing away $\lesssim 1$ fraction of $W$ in $\mathcal{W}$, we can assume that each $W$ in $\mathcal{W}$ satisfies $W\cap(\mathbb{F}_{p}\times\mathbf{0}^{k})=\mathbf{0}^{k+1}$. Denote $\widetilde{\mathbb{F}_{p}^{2}}=\mathbb{F}_{p}\times\mathbf{0}^{k-1}\times\mathbb{F}_{p}$. If $V\in\mathcal{V}_{W}$ and $W\in\mathcal{W}$, then $V\cap\widetilde{\mathbb{F}_{p}^{2}}$ is a line. We denote it by $\ell_{V}:=V\cap\widetilde{\mathbb{F}_{p}^{2}}$. Note that $\ell_{V}$ is not parallel to the first coordinate in $\widetilde{\mathbb{F}_{p}^{2}}$.
For a fixed $V\in\mathcal{V}_{W}$, we do a Fubini-type argument. Write
$V=\bigsqcup_{x\in\ell_{V}}(x+W).$
Noting that $Y(V)\sim p^{s}$, by pigeonholing, we find a subset $Y(\ell_{V})\subset\ell_{V}$, so that $Y(\ell_{V})\sim s_{1}(V)$, and for each $x\in Y(\ell_{V})$, $Y(V)\cap(x+W)\sim s_{2}(V)$. Here, $1\leq p^{s_{1}(V)}\leq p,1\leq p^{s_{2}(V)}\leq p^{s}$ are dyadic numbers that satisfy
$p^{s_{1}(V)}\cdot p^{s_{2}(V)}\gtrapprox p^{s}.$
By throwing away some horizontal slices of $Y(V)$, we can assume
$Y(V)=\bigsqcup_{x\in Y(\ell_{V})}\bigg(Y(V)\cap(x+W)\bigg).$
For the new $Y(V)$, we still have
$Y(V)\gtrapprox p^{s}.$
For convenience, given $V\in\mathcal{V}_{W}$ and $x\in\ell_{V}$, we denote
$Y(V,x):=Y(V)\cap(x+W).$
$Y(V,x)$ is called the horizontal $x$-slice. (Note that there is no confusion since $V$ uniquely determines $W$.) Thus, we have
$Y(V)=\bigsqcup_{x\in Y(\ell_{V})}Y(V,x).$ (33)
Third pigeonholing. For each $W\in\mathcal{W}$, we can find dyadic numbers $p^{s_{1}(W)},p^{s_{2}(W)}$ so that $\gtrapprox 1$ fraction of $V$ in $\mathcal{V}(W)$ satisfy $p^{s_{1}(V)}=p^{s_{1}(W)},p^{s_{2}(V)}=p^{s_{2}(W)}$.
Fourth pigeonholing. There exist dyadic numbers $p^{s_{1}},p^{s_{2}}$ such that $\gtrapprox 1$ fraction of $W$ in $\mathcal{W}$ satisfy $p^{s_{1}(W)}=p^{s_{1}}$ and $p^{s_{2}(W)}=p^{s_{2}}$.
We summarize what we have done so far. We can redefine $\mathcal{W}$, $\mathcal{V}_{W}$ and $\mathcal{V}$, so that $\mathcal{V}=\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$ and the following uniformity hold.
$\mathcal{W}\gtrapprox p^{t_{1}}.$
$\mathcal{V}_{W}\gtrapprox p^{t_{2}},\textup{~for~}W\in\mathcal{W}.$
For $W\in\mathcal{W}$ and each $V\in\mathcal{V}_{W}$,
$Y(V)=\bigsqcup_{x\in Y(\ell_{V})}Y(V,x),$ (34)
with $Y(\ell_{V})\sim p^{s_{1}}$ and for each $x\in Y(\ell_{V})$, $Y(V,x)\sim p^{s_{2}}$.
[…]
Lemma 4.5.
Assume Conjecture 1.5 holds. Suppose $E=\bigcup_{\ell\in\mathcal{L}}Y(\ell)$ is a $(s_{1},t_{2};2,1)$-set. In addition, we assume that each $\ell\in\mathcal{L}$ is not parallel to $\mathbb{F}_{p}\times\mathbf{0}$. Then there exist dyadic numbers $p^{u},p^{v}$ such that
$p^{s_{1}}\lesssim p^{u}\lesssim p,$
$1\leq p^{v}\leq p,$
and
$p^{u}p^{v}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
for any $\varepsilon>0$. In addition, $E$ contains a subset of the form
$\bigsqcup_{\xi\in\Xi}Y_{\xi},$
where $\Xi\subset\mathbf{0}\times\mathbb{F}_{p}$ with $\Xi\sim p^{u}$, and $Y_{\xi}\subset E\cap\bigg(\xi+(\mathbb{F}_{p}\times\mathbf{0})\bigg)$ with $Y_{\xi}\sim p^{v}$.
Proof.
For $\xi\in\mathbf{0}\times\mathbb{F}_{p}$, define $E(\xi):=E\cap\bigg(\xi+(\mathbb{F}_{p}\times\mathbf{0})\bigg)$. By pigeonholing on horizontal slices of $E$, we can find $\Xi\subset\mathbf{0}\times\mathbb{F}_{p}$ so that for all $\xi\in\Xi$, $E(\xi)$ have comparable cardinality. We may denote $\Xi\sim p^{u},\#E(\xi)\sim p^{v}$ for two dyadic numbers $p^{u},p^{v}\geq 1$. We can also assume $p^{u}p^{v}\gtrapprox\#E$. Hence, assuming Conjecture 1.5, we have
$p^{u}p^{v}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
However, we still need the condition $p^{u}\gtrsim p^{s_{1}}$. This is handled by the following trick. We will perform an algorithm to find a sequence $E=E_{1}\supset E_{2}\supset\cdots\supset E_{m}$, and disjoint subsets $\Xi_{1},\Xi_{2}\cdots,\Xi_{m}$ of $\mathbf{0}\times\mathbb{F}_{p}$. To start with, let $E_{1}=E$. By pigeonholing, we can find $\Xi_{1}\subset\mathbf{0}\times\mathbb{F}_{p}$, such that $\Xi_{1}\sim p^{u_{1}}$, $E(\xi)\sim p^{v_{1}}$ for each $\xi\in\Xi$, and
$p^{u_{1}}p^{v_{1}}\gtrapprox\#E_{1}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
Suppose that we have defined $E_{i},\Xi_{i}$ for $1\leq i\leq j-1$. If $\sum_{i=1}^{j-1}p^{u_{i}}\ll p^{s_{1}}$, then we define
$E_{j}:=E\setminus\bigcup_{1\leq i\leq j-1}\bigsqcup_{\xi\in\Xi_{i}}E(\xi).$
Since each $\ell\in\mathcal{L}$ is not parallel to $\mathbb{F}_{p}\times\mathbf{0}$, we see that $E_{j}$ is still a $(s_{1},t_{2};2,1)$-set. By the same pigeonholing, we find $\Xi_{j}\subset(\mathbf{0}\times\mathbb{F}_{p})\setminus\bigcup_{1\leq i\leq j-1}\Xi_{i}$, so that $\Xi_{j}\sim p^{u_{j}}$, $E(\xi)\sim p^{v_{j}}$ for each $\xi\in\Xi_{j}$, and
$p^{u_{j}}p^{v_{j}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
If $\sum_{i=1}^{j-1}p^{u_{i}}\gtrsim p^{s_{1}}$, then we stop and let m=j1m=j-1.
The algorithm will stop in finite steps. Denote $\Xi:=\bigcup_{1\leq i\leq m}\Xi_{i}$. Let $p^{u},p^{v}$ be dyadic numbers so that $p^{v}=\min_{1\leq i\leq m}p^{v_{i}}$, $p^{u}\sim\#\Xi$. For each $\xi\in\Xi$, let $Y_{\xi}$ be a subset of $E(\xi)$ with $Y_{\xi}\sim p^{v}$. This is doable, since $E(\xi)\gtrsim p^{v}$ if $\xi\in\Xi$. One can check
$p^{u}p^{v}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
This is proved in the following way. Suppose $p^{v}=p^{v_{i}}$, then
$p^{u}p^{v}=p^{u}p^{v_{i}}\geq p^{u_{i}}p^{v_{i}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};2,1)-\varepsilon}.$
Also, one can check
$p^{u}=\sum_{i=1}^{m}p^{u_{i}}\gtrsim p^{s_{1}}.$
Return to $\widetilde{E}_{W}:=\bigcup_{V\in\mathcal{V}_{W}}Y(\ell_{V})$ (which lies in $\widetilde{\mathbb{F}_{p}^{2}}=\mathbb{F}_{p}\times\mathbf{0}^{k-1}\times\mathbb{F}_{p}$). Recall $\mathcal{V}_{W}\gtrapprox p^{t_{2}}$ and $Y(\ell_{V})\sim p^{s_{1}}$. We see that $\widetilde{E}_{W}$ is a $(s_{1},(t_{2}-\varepsilon)_{+};2,1)$-set, for any $\varepsilon>0$.
By Lemma 4.5, there exist dyadic numbers $p^{u(W)},p^{v(W)}$, so that $\widetilde{E}_{W}$ contains a subset of form
$\bigsqcup_{\xi\in\Xi_{W}}Y_{W}(\rho_{\xi}).$
Here, $\Xi_{W}\subset\mathbf{0}^{k}\times\mathbb{F}_{p}$ with $\Xi_{W}\sim p^{u(W)}$; $\rho_{\xi}:=\xi+(\mathbb{F}_{p}\times\mathbf{0}^{k})$ and $Y_{W}(\rho_{\xi})\subset\rho_{\xi}$ with $Y_{W}(\rho_{\xi})\sim p^{v(W)}$. We also have
$p^{u(W)}\gtrsim p^{s_{1}}.$
$p^{u(W)}p^{v(W)}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},(t_{2}-\varepsilon)_{+};2,1)-\varepsilon}.$
By pigeonholing on $(u(W),v(W))$ again, we can throw away some $W\in\mathcal{W}$ so that the cardinality of $\mathcal{W}$ is still $\gtrapprox p^{t_{1}}$, and assume that there exist $u,v$ such that
$(u(W),v(W))=(u,v),\ \textup{~for~all~}W\in\mathcal{W},$
and
$p^{u}\gtrsim p^{s_{1}}.$
$p^{u}p^{v}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},(t_{2}-\varepsilon)_{+};2,1)-\varepsilon}.$
By Lemma 2.11,
$p^{u}p^{v}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
Finally, we are about to estimate the lower bound of the cardinality of $E=\bigcup_{W\in\mathcal{W}}\bigcup_{V\in\mathcal{V}_{W}}Y(V).$ This $E$ is not the original one, but the one after pigeonholing. Recall (34), we have
$\begin{split}\bigcup_{V\in\mathcal{V}_{W}}Y(V)&=\bigcup_{V\in\mathcal{V}_{W}}\bigsqcup_{x\in Y(\ell_{V})}Y(V,x)\\ &=\bigcup_{x\in\widetilde{E}_{W}}\bigg(\bigcup_{V:V\in\mathcal{V}_{W},x\in Y(\ell_{V})}Y(V,x)\bigg).\end{split}$ (35)
We define
$Y(W,x):=\bigcup_{V:V\in\mathcal{V}_{W},x\in Y(\ell_{V})}Y(V,x).$
Then $Y(W,x)$ is contained in the (k1)(k-1)-plane $x+W$.
If $x\in\widetilde{E}_{W}$, then
$Y(W,x)\gtrsim p^{s_{2}}.$
Recall that
$\widetilde{E}_{W}\supset\bigsqcup_{\xi\in\Xi_{W}}Y_{W}(\rho_{\xi}).$
We have
$\bigcup_{V\in\mathcal{V}_{W}}Y(V)\supset\bigcup_{\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x).$
The reader can check that, for fixed $\xi\in\Xi_{W}$, $\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x)$ lies in $\mathbb{F}_{p}^{k}\times\{\xi\}$.
Now we see that
$E\supset\bigcup_{W\in\mathcal{W}}\bigcup_{\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x).$
We want to estimate the lower bound of the right-hand side. Intuitively, one can imagine that the worst case is when $\Xi_{W}$ are morally the same for all $W\in\mathcal{W}$; otherwise we may translate them to the same $(k+1)$-th coordinate so that their union becomes smaller. For example, if we have sets $A_{i}\times\{\xi_{i}\}\subset\mathbb{F}_{p}^{k}\times\mathbb{F}_{p}$ ($i\in I$). Then $\bigcup_{i\in I}A_{i}\times\{\xi_{i}\}$ is the smallest when all the $\xi_{i}$ are the same.
Let us return to the estimate of
$E^{\prime}:=\bigcup_{W\in\mathcal{W}}\bigcup_{\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x).$
There are two lower bounds. The first lower bound is obtained from a fixed $W$. We have
$E^{\prime}\geq\bigcup_{\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x)\gtrsim p^{u}p^{v}p^{s_{2}}.$
For the second lower bound, we will estimate $E^{\prime}$ on each $\xi$-slice. We define the $\xi$-slice of $E^{\prime}$ by
$E^{\prime}_{\xi}:=E^{\prime}\cap(\mathbb{F}_{p}^{k}\times\{\xi\})=\bigcup_{W:W\in\mathcal{W},\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x).$
We will see that $E^{\prime}_{\xi}$ is some sort of $(s_{2},t_{\xi};k,k-1)$-set. Define
$p^{t_{\xi}}:=\#\{(W,x):W\in\mathcal{W},\xi\in\Xi_{W},x\in Y_{W}(\rho_{\xi})\}.$
We note two bounds for $t_{\xi}$:
$p^{t_{\xi}}\leq\#\mathcal{W}\cdot\#Y_{W}(\rho_{\xi})\lesssim p^{t_{1}}p^{v};$ (36)
$\sum_{\xi\in\mathbf{0}^{k}\times\mathbb{F}_{p}}p^{t_{\xi}}=\sum_{W\in\mathcal{W}}\#\Xi_{W}\cdot\#Y_{W}(\rho_{\xi})\gtrapprox p^{t_{1}}p^{u}p^{v}.$ (37)
We used that $p^{t_{1}}\gtrsim\#\mathcal{W}\gtrapprox p^{t_{1}}$, $\Xi_{W}\sim p^{u}$, $Y_{W}(\rho_{\xi})\sim p^{v}$.
And note that $Y(W,x)$ is a subset of the line $x+W$ such that $Y(W,x)\gtrsim p^{s_{2}}$. Therefore, $E_{\xi}^{\prime}$ is a $(s_{2},t_{\xi};k,k-1)$-set. By (27), we have
$E^{\prime}_{\xi}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}.$
Therefore,
$E\gtrsim_{\varepsilon}\sum_{\xi\in\mathbf{0}^{k}\times\mathbb{F}_{p}}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}.$
We claim that the right-hand side is
$\gtrapprox p^{u}p^{\mathbf{F}(s_{2},t_{1}+v;k,k-1)-\varepsilon}.$
By pigeonholing, we can find a dyadic number $p^{t_{\circ}}\lesssim p^{t_{1}}p^{v}$ and $\Xi_{\circ}\subset\mathbf{0}^{k}\times\mathbb{F}_{p}$, so that $2p^{t_{\circ}}\geq p^{t_{\xi}}\geq p^{t_{\circ}}$ for $\xi\in\Xi_{\circ}$ and
$\Xi_{\circ}\cdot p^{t_{\circ}}\gtrapprox p^{t_{1}}p^{u}p^{v}.$
Hence,
$\Xi_{\circ}\gtrapprox p^{u}p^{t_{1}+v-t_{\circ}}.$
We can estimate
$E\gtrsim_{\varepsilon}\sum_{\xi\in\Xi_{\circ}}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}\geq\#\Xi_{\circ}\cdot p^{\mathbf{F}(s_{2},t_{\circ};k,k-1)-\varepsilon}\gtrapprox p^{u}p^{\mathbf{F}(s_{2},t_{\circ};k,k-1)+t_{1}+v-t_{\circ}-\varepsilon}.$
By Lemma 2.11, we have
$E\gtrapprox_{\varepsilon}p^{u+\mathbf{F}(s_{2},t_{1}+v;k,k-1)-\varepsilon}.$
Note that the implicit constants in the above arguments may depend on various parameters $s_{1},s_{2},t_{1},t_{2}$ and so on. To get rid of this dependence, we use the trick discussed before this lemma. Hence, we update the estimate to
$p^{u+v}\gtrsim_{\varepsilon}p^{\mathbf{F}((s_{1}-\varepsilon^{2})_{+},(t_{2}-\varepsilon^{2})_{+};2,1)-\varepsilon},$
$E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}((s_{2}-\varepsilon^{2})_{+},(t_{1}+v-\varepsilon^{2})_{+};k,k-1),s_{2}+v\}-\varepsilon}.$
Now, the implicit constants only depend on $s,t,k,\varepsilon$. ∎
The next lemma is a recursive inequality of $\mathbf{F}(s,t;k+1,k)$, which can be used to bound the right-hand side of (31).
Lemma 4.6.
Suppose $(s,t;k+1,k)$ is admissible. Suppose there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1},1],v\in[0,1];\end{cases}$ (38)
and
$\begin{cases}t_{1}+t_{2}\geq t;\\ s_{1}+s_{2}\geq s;\\ u+v\geq\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$ (39)
Then,
$u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$ (40)
Let us quickly see how Lemma 4.4 and Lemma 4.6 combined to prove Proposition 4.3.
Proof of Proposition 4.3 assuming Lemma 4.4 and Lemma 4.6.
If $s=0$ or $t=0$, by Proposition 4.1 or 4.2, we are done. So, we suppose $s,t>0$.
Fix $0<\varepsilon<\min\{\sigma/2,t/2,1/2\}$, where $s=d+\sigma$ with $d\in\mathbb{N},\sigma\in(0,1]$. We just need to prove
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k+1,k)-C\varepsilon}$ (41)
for any $(s,t;k+1,k)$-set $E$. Here, $C$ is a large number.
We only need to prove for large enough $p$, since for small $p$ we can choose the constant to be large.
Suppose (27) holds. Then in particular, Conjecture 1.5 also holds. If $E$ is a $(s,t;k+1,k)$-set, by Lemma 4.4, there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1}-\varepsilon,1],v\in[0,1];\end{cases}$ (42)
and
$\begin{cases}t_{1}+t_{2}\geq t-\varepsilon;\\ s_{1}+s_{2}\geq s-\varepsilon;\\ u+v\geq\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon;\end{cases}$ (43)
and
$E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}.$ (44)
Here, we renamed $t_{1}$ to be $(t_{1}-\varepsilon^{2})_{+}$ in Lemma 4.4 (and similar for $t_{2},s_{1},s_{2},u,v$). Also, we used the fact that when $p$ is large enough (depending on $\varepsilon$), then $p^{t_{1}+t_{2}}\gtrapprox p^{t}$ implies $t_{1}+t_{2}\geq t-\varepsilon$. Also, by our assumption. $t-\varepsilon,s-\varepsilon>0$. Let $t^{\prime}=t-\varepsilon$. $s^{\prime}=s-\varepsilon$.
Since $\mathbf{F}(s_{1},t_{2};2,1)\leq 2$, we can find $u\leq u^{\prime}\leq\min\{u+\varepsilon,1\}$ and $v\leq v^{\prime}\leq\min\{v+\varepsilon,1\}$, so that $u^{\prime}\geq s_{1}$, $u^{\prime}+v^{\prime}\geq\mathbf{F}(s_{1},t_{2};2,1)$. Now we have
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u^{\prime}\in[s_{1},1],v^{\prime}\in[0,1];\end{cases}$ (45)
and
$\begin{cases}t_{1}+t_{2}\geq t^{\prime};\\ s_{1}+s_{2}\geq s^{\prime};\\ u^{\prime}+v^{\prime}\geq\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$ (46)
and
$E\gtrsim p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}\geq p^{u^{\prime}+\max\{\mathbf{F}(s_{2},t_{1}+v^{\prime};k,k-1),s_{2}+v^{\prime}\}-O(\varepsilon)}.$ (47)
In the last inequality, we used Lemma 2.11.
Finally, by Lemma 4.6, we have
$u^{\prime}+\max\{\mathbf{F}(s_{2},t_{1}+v^{\prime};k,k-1),s_{2}+v^{\prime}\}\geq\mathbf{F}(s^{\prime},t^{\prime};k+1,k)\geq\mathbf{F}(s,t;n,k)-O(\varepsilon).$
The last inequality is by Proposition 2.12 and 2.11.
In summary, we finish the proof of
$E\gtrsim p^{\mathbf{F}(s,t;k+1,k)-C\varepsilon}.$
It remains to prove Lemma 4.6. It can be reduced to the following lemma, with $\geq$” in (39) all replaced by “==”.
Lemma 4.7.
Suppose $(s,t;k+1,k)$ is admissible. Suppose there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
$\begin{cases}t_{1}\in[0,k-1],t_{2}\in[0,2];\\ s_{1}\in[0,1],s_{2}\in[0,s];\\ u\in[s_{1},1],v\in[0,1];\end{cases}$ (48)
and
$\begin{cases}t_{1}+t_{2}=t;\\ s_{1}+s_{2}=s;\\ u+v=\mathbf{F}(s_{1},t_{2};2,1);\end{cases}$ (49)
Then,
$u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$ (50)
We show that Lemma 4.7 implies Lemma 4.6.
Proof of Lemma 4.6 assuming Lemma 4.7.
The idea is to slightly modify $s_{1},s_{2},t_{1},t_{2},u,v$ so that the $\geq$” in (39) becomes “==”, and then we can apply Lemma 4.7.
For $i=1,2$, we can find $t_{i}^{\prime}\in[0,t_{i}],s^{\prime}_{i}\in[0,s_{i}]$, so that $t_{1}^{\prime}+t_{2}^{\prime}=t,s_{1}^{\prime}+s_{2}^{\prime}=s$. It is slightly trickier to modify $u,v$. Noting that $u+v\geq\mathbf{F}(s_{1},t_{2};2,1)\geq\mathbf{F}(s_{1}^{\prime},t_{2}^{\prime};2,1)\geq s_{1}^{\prime}$, we can find $u^{\prime}\in[s_{1}^{\prime},u],v^{\prime}\in[0,v]$ so that $u^{\prime}+v^{\prime}=\mathbf{F}(s_{1}^{\prime},t_{2}^{\prime};2,1)$. Now we can apply Lemma 4.7 to the numbers $t_{1}^{\prime},t_{2}^{\prime},s_{1}^{\prime},s_{2}^{\prime},u^{\prime},v^{\prime}$. We obtain
$u^{\prime}+\max\{\mathbf{F}(s^{\prime}_{2},t^{\prime}_{1}+v^{\prime};k,k-1),s^{\prime}_{2}+v^{\prime}\}\geq\mathbf{F}(s,t;k+1,k).$
Since the new numbers are smaller than the old numbers and by using the monotonicity of $\mathbf{F}(\cdot,\cdot;n,k)$, we have
$u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}\geq\mathbf{F}(s,t;k+1,k).$
It remains to prove Lemma 4.7.
Proof of Lemma 4.7.
For convenience, we recall the definition of Furstenberg index in the special case $(n,k)=(k+1,k)$.
(a) If $s=0$,
$\mathbf{F}(0,t;k+1,k)=\max\{0,t-k\}.$
(b) If $s>0$, write $s=d+\sigma$. If also $t\in[0,k-d-1]$,
$\mathbf{F}(s,t;k+1,k)=s.$
If we are not in Case (a) and (b), then $s\in(0,k]$ and $t\in(k-d-1,k+1]$, we write
$\begin{cases}s=d+\sigma\\ t=k-d-1+\tau,\end{cases}$
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $\tau\in(0,d+2]$.
(c) If $\tau\in(0,2]$,
$\mathbf{F}(s,t;k+1,k)=d+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}$
(d) If $\tau\in(2,d+2]$,
$\mathbf{F}(s,t;k+1,k)=s+1.$
Let us denote
$\mathbf{F}:=u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}.$
Our goal is to prove
$\mathbf{F}\geq\mathbf{F}(s,t;k+1,k).$
When $s=0$, then $s_{1}=s_{2}=0$. We want to prove $\mathbf{F}\geq\max\{0,t-k\}$. Trivially, $\mathbf{F}\geq 0$. Also,
$\mathbf{F}\geq u+\mathbf{F}(0,t_{1}+v;k,k-1)\geq u+t_{1}+v-(k-1).$
Using $u+v=\mathbf{F}(s_{1},t_{2};2,1)\geq t_{2}-1$ and $t_{1}+t_{2}=t$, we obtain
$\mathbf{F}\geq t-k.$
When $t\in[0,k-d-1]$, we have
$\mathbf{F}\geq u+s_{2}+v=s_{2}+\mathbf{F}(s_{1},t_{2};2,1)\geq s_{2}+s_{1}=s=\mathbf{F}(s,t;k+1,k).$
So we consider $s>0$ and t>kd1t>k-d-1. Write
$\begin{cases}s=d+\sigma\\ t=k-d-1+\tau,\end{cases}$
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $\tau\in(0,d+2]$. In this case,
$\mathbf{F}(s,t;k+1,k)=d+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}.$
We just need to show either $\mathbf{F}\geq d+\sigma+\tau$, or $\mathbf{F}\geq d+\frac{3}{2}\sigma+\frac{1}{2}\tau$, or $\mathbf{F}\geq d+\sigma+1$.
We first analyze $\mathbf{F}(s_{2},t_{1}+v;k,k-1)$. This is done by discussing the values of $(s_{2},t_{1}+v)$.
Case (a): $s_{2}=0$By definition, we have $\mathbf{F}(s_{2},t_{1}+v;k,k-1)=\max\{0,t_{1}+v-(k-1)\}$. Since $s_{1}+s_{2}=s=d+\sigma$ and $s_{1}\leq 1$, we have $d=0,s_{1}=\sigma$. We have
$\mathbf{F}\geq u+\max\{t_{1}+v-(k-1),v\}.$
Recall $u+v\geq\mathbf{F}(\sigma,t_{2};2,1),t_{1}+t_{2}=t=k-1+\tau$. We have
$\mathbf{F}\geq\mathbf{F}(\sigma,t_{2};2,1)+\max\{\tau-t_{2},0\}.$
If $t_{2}\geq\tau$, then $\mathbf{F}(\sigma,t_{2};2,1)\geq\mathbf{F}(\sigma,\tau;2,1)$. If $t_{2}\leq\tau$, then $\mathbf{F}(\sigma,t_{2};2,1)+\tau-t_{2}\geq\mathbf{F}(\sigma,\tau;2,1)$, by Lemma 2.11. Therefore, we have
$\mathbf{F}\geq\mathbf{F}(\sigma,\tau;2,1)=\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}=\mathbf{F}(s,t;k+1,k).$
The last inequality is because $d=0$.
Case (b) $t_{1}+v\leq k-d-2$ We have
$\mathbf{F}\geq u+v+s_{2}\geq\mathbf{F}(s_{1},t_{2};2,1)+s_{2}.$
Subcase (b.0): $s_{1}=0$In this case, $\mathbf{F}(s_{1},t_{2};2,1)=\max\{0,t_{2}-1\}$, $s_{2}=d+\sigma$. We have
$\mathbf{F}\geq\max\{0,t_{2}-1\}+d+\sigma\geq d+\sigma+\tau.$
The last inequality is because
$t_{2}=t-t_{1}\geq(k-d-1+\tau)-(k-d-2-v)\geq 1+\tau.$
Subcase (b.1): $\mathbf{F}(s_{1},t_{2};2,1)=s_{1}+t_{2}$We have
$\mathbf{F}\geq s_{1}+t_{2}+s_{2}=s+t_{2}\geq d+\sigma+\tau.$
since $t_{2}\geq 1+\tau$.
Subcase (b.2): $\mathbf{F}(s_{1},t_{2};2,1)=\frac{3}{2}s_{1}+\frac{1}{2}t_{2}$We have
$\mathbf{F}\geq\frac{1}{2}(s_{1}+t_{2})+d+\sigma\geq d+\frac{3}{2}\sigma+\frac{1}{2}\tau.$
Since $t_{2}\geq 1+\tau\geq\sigma+\tau.$
Subcase (b.3): $\mathbf{F}(s_{1},t_{2};2,1)=s_{1}+1$We have
$\mathbf{F}\geq s+1=d+\sigma+1.$
Case (c): $s_{2}>0$, $k-d-2<t_{1}+v\leq k-d$We write $t_{1}+v=k-d-2+\tau^{\prime}$ for $\tau^{\prime}\in(0,2]$. Since $t_{1}+t_{2}=k-d-1+\tau$, we have $\tau^{\prime}=1+\tau+v-t_{2}$.
Subcase (c.I): $s_{1}<\sigma$ We have $s_{2}=s-s_{1}=d+(\sigma-s_{1})$. We have
$\begin{split}\mathbf{F}(s_{2},t_{1}+v;k,k-1)=\mathbf{F}(d+(\sigma-s_{1}),t_{1}+v;k,k-1)\\ =d+\min\{\sigma-s_{1}+\tau^{\prime},\frac{3}{2}(\sigma-s_{1})+\frac{1}{2}\tau^{\prime},(\sigma-s_{1})+1\}.\end{split}$ (51)
Subsubcase (c.I.1): $\mathbf{F}(s_{2},t_{1}+v;k,k-1)=s_{2}+\tau^{\prime}$We have
$\begin{split}\mathbf{F}&\geq u+\max\{s_{2}+\tau^{\prime},s_{2}+v\}\\ &=u+s_{2}+\max\{\tau^{\prime},v\}.\end{split}$ (52)
Subsubsubcase (c.I.1.0): $s_{1}=0$Then $u+v=\mathbf{F}(0,t_{2};2,1)=\max\{0,t_{2}-1\}$. We have
$\mathbf{F}\geq u+s_{2}+\tau^{\prime}=u+v+s_{2}+1+\tau-t_{2}\geq s_{2}+\tau=s+\tau.$
Subsubsubcase (c.I.1.1): $u+v=\mathbf{F}(s_{1},t_{2};2,1)=s_{1}+t_{2}$We have
$\mathbf{F}\geq u+s_{2}+\tau^{\prime}=u+s_{2}+1+\tau+v-t_{2}=s+\tau+1.$
Subsubsubcase (c.I.1.2): $u+v=\mathbf{F}(s_{1},t_{2};2,1)=\frac{3}{2}s_{1}+\frac{1}{2}t_{2}$We have
$\mathbf{F}\geq\frac{3}{2}s_{1}+\frac{1}{2}t_{2}+s_{2}+1+\tau-t_{2}=s+\frac{1}{2}(s_{1}+t_{2})+1+\tau-t_{2}\geq s+1+\tau-\frac{1}{2}t_{2}.$
We also have
$\mathbf{F}\geq u+v+s_{2}\geq s+\frac{1}{2}t_{2}.$
Combining together gives
$\mathbf{F}\geq s+\frac{1}{2}(1+\tau)\geq d+\frac{3}{2}\sigma+\frac{1}{2}\tau.$
Subsubsubcase (c.I.1.3): $u+v=\mathbf{F}(s_{1},t_{2};2,1)=s_{1}+1$We have
$\mathbf{F}\geq u+v+s_{2}\geq s+1.$
SubSubcase (c.I.2): $\mathbf{F}(s_{2},t_{1}+v;k,k-1)=d+\frac{3}{2}(\sigma-s_{1})+\frac{1}{2}\tau^{\prime}=s_{2}+\frac{1}{2}(\sigma-s_{1}+\tau^{\prime})$We have
$\mathbf{F}\geq u+s_{2}+\max\{\frac{1}{2}(\sigma-s_{1}+\tau^{\prime}),v\}=u+s_{2}+\max\{\frac{1}{2}(\sigma+1+\tau+v-t_{2}-s_{1}),v\}.$
Subsubsubcase (c,I.2.0): $s_{1}=0$Then $u+v=\max\{0,t_{2}-1\}$. We have
$\mathbf{F}\geq\frac{1}{2}u+s_{2}+\frac{1}{2}(u+v+\sigma+1+\tau-t_{2}-s_{1})\geq s_{2}+\frac{1}{2}(\sigma+\tau-s_{1})=d+\frac{3}{2}\sigma+\frac{1}{2}\tau.$
Subsubsubcase (c.I.2.1): $u+v=s_{1}+t_{2}$We have
$\mathbf{F}\geq\frac{1}{2}u+s_{2}+\frac{1}{2}(\sigma+\tau+1)\geq d+\frac{3}{2}\sigma+\frac{1}{2}\tau.$
Here, we used $u\geq s_{1},1\geq s_{1}$.
Subsubsubcase (c.I.2.2): $u+v=\frac{3}{2}s_{1}+\frac{1}{2}t_{2}$We have
$\mathbf{F}\geq\frac{1}{2}u+s_{2}+\frac{1}{2}(\sigma+1+\tau)+\frac{1}{4}(s_{1}-t_{2})$
To prove $\mathbf{F}\geq d+\frac{3}{2}\sigma+\frac{1}{2}\tau$, it suffices to prove
$\frac{1}{2}(u+1)+\frac{1}{4}(s_{1}-t_{2})\geq s_{1}.$
This is true since $u+1\geq u+v=\frac{3}{2}s_{1}+\frac{1}{2}t_{2}$.
Subsubsubcase (c.I.2.3): $u+v=s_{1}+1$
$\mathbf{F}\geq u+v+s_{2}=s+1.$
Subsubcase (c.I.3): $\mathbf{F}(s_{2},t_{1}+v;k,k-1)=s_{2}+1$ We have
$\mathbf{F}\geq u+s_{2}+1\geq s_{1}+s_{2}+1=s+1.$
Subcase (c.II): $s_{1}\geq\sigma$ We write $s_{2}=(d-1)+(\sigma+1-s_{1})$. We have
$\mathbf{F}(s_{2},t_{1}+v;k,k-1)$ $=d-1+\min\{\sigma+1-s_{1}+\tau^{\prime},\frac{3}{2}(\sigma-s_{1}+1)+\frac{1}{2}\tau^{\prime},\sigma-s_{1}+2\}$
$=d+\min\{\sigma-s_{1}+\tau^{\prime},\frac{3}{2}(\sigma-s_{1})+\frac{1}{2}\tau^{\prime}+\frac{1}{2},\sigma-s_{1}+1\}$
$\geq d+\min\{\sigma-s_{1}+\tau^{\prime},\frac{3}{2}(\sigma-s_{1})+\frac{1}{2}\tau^{\prime},\sigma-s_{1}+1\}$
The discussion is the same as Subcase (c.I), since we did not use the condition $s_{1}<\sigma$.
Subcase (d): $t_{1}+v>k-d$We have
$\mathbf{F}(s_{2},t_{1}+v;k,k-1)=s_{2}+1.$
Therefore,
$\mathbf{F}\geq u+s_{2}+1\geq s_{1}+s_{2}+1=s+1.$
5 Furstenberg set problem: Lower bound II
Proposition 5.1.
Let $n,k\in\mathbb{N}$ and $n\geq k+2$. Suppose for any $n^{\prime}=k+1,\dots,n-1$ and admissible $(s,t;n^{\prime},k)$, we know
$\inf_{E:\ (s,t;n^{\prime},k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n^{\prime},k)-\varepsilon},$ (53)
for any $\varepsilon>0$. Then for admissible $(s,t;n,k)$, we know
$\inf_{E:\ (s,t;n,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon},$ (54)
for any $\varepsilon>0$.
We prove the following lemma. Actually, we also need to deal with a technical thing as mentioned before Lemma 4.4. Since we have done it once in the proof of Lemma 4.4, we will not reproduce here.
Lemma 5.2.
Assume (53) holds. Suppose $(s,t;n,k)$ is admissible and $n\geq k+2$. Then for any $(s,t;n,k)$-set $E$, there exist numbers $t_{1},t_{2},s_{1},s_{2}$, such that
$\begin{cases}1\leq p^{t_{1}}\leq p^{(k+1)(n-k-1)},1\leq p^{t_{2}}\leq p^{k+1};\\ p^{s}\lesssim p^{s_{1}}\leq p^{k},1\leq p^{s_{2}}\leq p;\end{cases}$ (55)
and
$\begin{cases}p^{t_{1}+t_{2}}\gtrapprox p^{t};\\ p^{s_{1}+s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};n-1,k)-\varepsilon};\end{cases}$ (56)
and
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)+s_{2}-\varepsilon}.$ (57)
Proof.
We will do several steps of pigeonholing, which lead to the parameters $t_{1},t_{2},s_{1},s_{2},u,v$.
By the definition of $(s,t;n,k)$-set, we write
$E=\bigcup_{V\in\mathcal{V}}Y(V),$
where $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{n})$ with $\mathcal{V}\sim p^{t}$, and $Y(V)\subset V$ with $Y(V)\sim p^{s}$.
The strategy is to find a subset of $\mathcal{V}$, and a subset of $Y(V)$ for each $V$, still denoted by $\mathcal{V}$, $Y(V)$, so that the new $\mathcal{V},Y(V)$ satisfy certain uniformity. This is done by a sequence of pigeonhole arguments.
By properly choosing the coordinate, we can assume that $\gtrsim 1$ fraction of the $k$-planes in $\mathcal{V}$ are not parallel to the $n$-th axis $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$. Throwing away some $V\in\mathcal{V}$, we can just assume all the $k$-planes in $\mathcal{V}$ are not parallel to $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$ with $\mathcal{V}$ still satisfying $\mathcal{V}\gtrsim p^{t}$. Because of this, we see that for any $V\in\mathcal{V}$, the projection of $V$ onto $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$, denoted by $\pi_{\mathbb{F}_{p}^{n-1}}(V)$ is a $k$-plane in $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$.
Now, we partition $\mathcal{V}$ according to $A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$. For each $W\in A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$, we define $\mathcal{V}_{W}$ to be the set of $V\in\mathcal{V}$ such that $\pi_{\mathbb{F}_{p}^{n-1}}(V)=W$. We obtain a partition
$\mathcal{V}=\bigsqcup_{W\in A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})}\mathcal{V}_{W}.$
First pigeonholing. We assume that there exists $\mathcal{W}\subset A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with $\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, $\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $1\leq p^{t_{1}}\leq p^{(k+1)(n-k-1)},1\leq p^{t_{2}}\leq p^{k+1}$ are dyadic numbers, and
$p^{t_{1}}p^{t_{2}}\gtrapprox p^{t}.$ (58)
The upper bound for $t_{1},t_{2}$ is because $A(k,\mathbb{F}_{p}^{n-1}\times\mathbf{0})\lesssim p^{(k+1)(n-k-1)}$, and $\mathcal{V}_{W}\leq\#A(k,W\times\mathbb{F}_{p})\lesssim p^{k+1}$. We still use $\mathcal{V}$ to denote $\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$. And we see that
$\mathcal{V}\gtrapprox p^{t}.$
Fix a $W\in\mathcal{W}$. We want to estimate $\bigcup_{V\in\mathcal{V}_{W}}Y(V)(\subset W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{k+1})$. This is some sort of Furstenberg set for $k$-planes in $(k+1)$-dimensional space. Also, an important property is that each $V\in\mathcal{V}_{W}$ is not parallel to the last coordinate. Therefore, any line parallel to $\mathbf{0}^{n-1}\times\mathbb{F}_{p}$ will intersect $V$ at a single point. We will exploit some “quasi-product” structure of this set.
Lemma 5.3.
Assume (53) holds for $n^{\prime}=k+1$. Suppose $E=\bigcup_{V\in\mathcal{V}}Y(V)$ is a $(s,t_{2};k,k+1)$-set. In addition, we assume that each $V\in\mathcal{V}$ intersects $\mathbf{0}^{k}\times\mathbb{F}_{p}$ at a single point. Then there exist dyadic numbers $p^{s_{1}},p^{s_{2}}$ such that
$p^{s}\lesssim p^{s_{1}}\leq p^{k},$
$1\leq p^{s_{2}}\leq p,$
and
$p^{s_{1}}p^{s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};k+1,k)-\varepsilon},$
for any $\varepsilon>0$. In addition, $E$ contains a subset of form
$\bigsqcup_{\xi\in\Xi}Y_{\xi},$
where $\Xi\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$ with $\Xi\sim p^{s_{1}}$, and $Y_{\xi}\subset E\cap\bigg(\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})\bigg)$ with $Y_{\xi}\sim p^{s_{2}}$.
Proof.
For $\xi\in\mathbb{F}_{p}^{k}\times\mathbf{0}$, define $E(\xi):=E\cap\bigg(\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})\bigg)$. By pigeonholing on vertical slices of $E$, we can find $\Xi\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$ so that for all $\xi\in\Xi$, $E(\xi)$ have comparable cardinality. We may denote $\Xi\sim p^{s_{1}},\#E(\xi)\sim p^{s_{2}}$ for two dyadic numbers $p^{s_{1}},p^{s_{2}}\geq 1$. We can also assume $p^{s_{1}}p^{s_{2}}\gtrapprox\#E$. Hence, assuming (53) holds for $n^{\prime}=k+1$, we have
$p^{s_{1}}p^{s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};k+1,k)-\varepsilon}.$
However, we still need the condition $p^{s_{1}}\gtrsim p^{s}$. This is handled by the same trick as in the proof of Lemma 4.5. We will perform an algorithm to find a sequence $E=E_{1}\supset E_{2}\supset\cdots\supset E_{m}$, and disjoint subsets $\Xi_{1},\Xi_{2}\cdots,\Xi_{m}$ of $\mathbb{F}_{p}^{k}\times\mathbf{0}$. To start with, let $E_{1}=E$. By pigeonholing, we can find $\Xi_{1}\subset\mathbb{F}_{p}^{k}\times\mathbf{0}$, such that $\Xi_{1}\sim p^{u_{1}}$, $E(\xi)\sim p^{v_{1}}$ for each $\xi\in\Xi$, and
$p^{u_{1}}p^{v_{1}}\gtrsim_{\varepsilon}\#E_{1}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};2,1)-\varepsilon}.$
Suppose we have defined $E_{i},\Xi_{i}$ for $1\leq i\leq j-1$. If $\sum_{i=1}^{j-1}p^{u_{i}}\ll p^{s}$, then we define
$E_{j}:=E\setminus\bigcup_{1\leq i\leq j-1}\bigsqcup_{\xi\in\Xi_{i}}E(\xi).$
Since each $V\in\mathcal{V}$ intersects $\xi+(\mathbf{0}^{k}\times\mathbb{F}_{p})$ (and hence $E(\xi)$) at a single point, we see that $E_{j}$ is still a $(s,t_{2};2,1)$-set. By the same pigeonholing, we find $\Xi_{j}\subset(\mathbb{F}_{p}^{k}\times\mathbf{0})\setminus\bigcup_{1\leq i\leq j-1}\Xi_{i}$, so that $\Xi_{j}\sim p^{u_{j}}$, $E(\xi)\sim p^{v_{j}}$ for each $\xi\in\Xi_{j}$, and
$p^{u_{j}}p^{v_{j}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};2,1)-\varepsilon}.$
If $\sum_{i=1}^{j-1}p^{u_{i}}\gtrsim p^{s}$, then we stop and let m=j1m=j-1.
The algorithm will stop in finite steps. Denote $\Xi:=\bigcup_{1\leq i\leq m}\Xi_{i}$. Let $p^{s_{1}},p^{s_{2}}$ be dyadic numbers so that $p^{s_{2}}=\min_{1\leq i\leq m}p^{v_{i}}$, $p^{s_{1}}\sim\#\Xi$. For each $\xi\in\Xi$, let $Y_{\xi}$ be a subset of $E(\xi)$ with $Y_{\xi}\sim p^{s_{2}}$. This is doable, since $E(\xi)\gtrsim p^{s_{2}}$ if $\xi\in\Xi$. One can check
$p^{s_{1}}p^{s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};2,1)-\varepsilon}.$
This is proved in the following way. Suppose $p^{s_{2}}=p^{v_{i}}$, then
$p^{s_{1}}p^{s_{2}}=p^{s_{1}}p^{v_{i}}\geq p^{u_{i}}p^{v_{i}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{1};2,1)-\varepsilon}.$
Also, one can check
$p^{s_{1}}=\sum_{i=1}^{m}p^{u_{i}}\gtrsim p^{s}.$
Return to $E_{W}:=\bigcup_{V\in\mathcal{V}_{W}}Y(V)$ (which lies in $W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{k+1}$). Recall $\mathcal{V}_{W}\gtrapprox p^{t_{2}}$ and $Y(V)\sim p^{s}$. We see that $E_{W}$ is a $(s,(t_{2}-\varepsilon)_{+};k+1,k)$-set.
By Lemma 5.3, there exist dyadic numbers $p^{s_{1}(W)},p^{s_{2}(W)}$, so that $E_{W}$ contains a subset of form
$\bigsqcup_{\xi\in\Xi_{W}}Y_{W}(\rho_{\xi}).$
Here, $\Xi_{W}\subset W(\subset\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with $\Xi_{W}\sim p^{s_{1}(W)}$; $\rho_{\xi}:=\xi+(\mathbf{0}^{n-1}\times\mathbb{F}_{p})$ and $Y_{W}(\rho_{\xi})\subset\rho_{\xi}$ with $Y_{W}(\rho_{\xi})\sim p^{s_{2}(W)}$. We also have
$p^{s_{1}(W)}\gtrsim p^{s}.$
$p^{s_{1}(W)}p^{s_{2}(W)}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,(t_{2}-\varepsilon)_{+};k+1,k)-\varepsilon}.$
By pigeonholing on $(s_{1}(W),s_{2}(W))$ again, we can throw away some $W\in\mathcal{W}$ so that the cardinality of $\mathcal{W}$ is still $\gtrapprox p^{t_{1}}$, and assume that there exist $s_{1},s_{2}$ such that
$(s_{1}(W),s_{2}(W))=(s_{1},s_{2}),\ \textup{~for~all~}W\in\mathcal{W},$
and
$p^{s_{1}}\gtrsim p^{s}.$
$p^{s_{1}}p^{s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,(t_{2}-\varepsilon)_{+};2,1)-\varepsilon}.$
By Lemma 2.11, we have
$p^{s_{1}}p^{s_{2}}\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t_{2};2,1)-\varepsilon}.$
Finally, we are about to estimate the lower bound of the cardinality of $E=\bigcup_{W\in\mathcal{W}}\bigcup_{V\in\mathcal{V}_{W}}Y(V).$ This $E$ is not the original one, but the one after pigeonholing. We write
$\begin{split}E=\bigcup_{W\in\mathcal{W}}E_{W}&\supset\bigcup_{W\in\mathcal{W}}\bigcup_{\xi\in\Xi_{W}}Y_{W}(\rho_{\xi}).\end{split}$ (59)
Since $Y_{W}(\rho_{\xi})\sim p^{s_{2}}$, we see that
$E\gtrsim\#(\bigcup_{W\in\mathcal{W}}\Xi_{W})\cdot p^{s_{2}}.$
Since $\bigcup_{W\in\mathcal{W}}\Xi_{W}$ is a $(s_{1},t_{1};n-1,k)$-set, by (53), we have
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)-\varepsilon}p^{s_{2}}.$
The next lemma is a recursive inequality of $\mathbf{F}(s,t;n,k)$, which can be used to bound the right-hand side of (57).
Lemma 5.4.
Suppose $(s,t;n,k)$ is admissible and $n\geq k+2$. Suppose that there exist numbers $t_{1},t_{2},s_{1}$, such that
$\begin{cases}t_{1}\in[0,(k+1)(n-k-1)];\\ t_{2}\in[0,k+1];\\ s_{1}\in[s,k];\end{cases}$ (60)
and
$t_{1}+t_{2}=t;$ (61)
Then,
$\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}\geq\mathbf{F}(s,t;n,k).$ (62)
Let us quickly see how Lemma 5.2 and Lemma 5.4 combined to prove Proposition 5.1.
Proof of Proposition 5.1 assuming Lemma 5.2 and Lemma 5.4.
If $s=0$ or $t=0$, by Proposition 4.1 or 4.2, we are done. So, we suppose $s,t>0$.
Fix $0<\varepsilon<\min\{\sigma/2,t/2,1/2\}$, where $s=d+\sigma$ with $d\in\mathbb{N},\sigma\in(0,1]$. We just need to prove
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-C\varepsilon}$ (63)
for any $(s,t;n,k)$-set $E$. Here, $C$ is a large number.
We only need to prove for large enough $p$, since for small $p$ we can choose the constant to be large.
Suppose (53) holds. If $E$ is a $(s,t;n,k)$-set, by Lemma 5.2, there exist numbers $t_{1},t_{2},s_{1},s_{2}$, such that
$\begin{cases}t_{1}\in[0,(k+1)(n-k-1)],t_{2}\in[0,k+1];\\ s_{1}\in[s-\varepsilon,k],s_{2}\in[0,1];\end{cases}$ (64)
and
$\begin{cases}t_{1}+t_{2}\geq t-\varepsilon;\\ s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};k+1,k)-\varepsilon;\end{cases}$ (65)
and
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)+s_{2}-\varepsilon}\geq p^{\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}-2\varepsilon}.$ (66)
Let $t^{\prime}=t-\varepsilon$ which is >0>0 by assumption. Choose $0\leq t_{i}^{\prime}\leq t_{i}$ for $i=1,2$, so that $t_{1}^{\prime}+t_{2}^{\prime}=t^{\prime}$. Choose $s_{1}\leq s_{1}^{\prime}\leq s_{1}+\varepsilon$, so that $s_{1}^{\prime}\in[s,k]$. Then $t^{\prime},t_{1}^{\prime},t_{2}^{\prime},s_{1}^{\prime}$ satisfy the condition in Lemma 5.4, we obtain
$\mathbf{F}(s_{1}^{\prime},t_{1}^{\prime};n-1,k)+\max\{\mathbf{F}(s,t_{2}^{\prime};k+1,k)-s_{1},0\}\geq\mathbf{F}(s,t^{\prime};n,k).$
By the monotonicity of $t\mapsto\mathbf{F}(s,t;n,k)$, Lemma 2.11 and Lemma 2.12, we have
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-C\varepsilon}.$
It remains to prove Lemma 5.4.
Proof of Lemma 5.4.
When $s=0$, we have $\mathbf{F}(0,t;n,k)=\max\{0,t-k(n-k)\}$. By Lemma 2.10, we have
$\mathbf{F}(s_{1},t_{1};n-1,k)+\mathbf{F}(0,t_{2};k+1,k)-s_{1}\geq s_{1}+t_{1}-k(n-k-1)+t_{2}-k-s_{1}=t-k(n-k).$
When $s>0$, write $s=d+\sigma$. If also $t\leq(k-d-1)(n-k)$, then
$\mathbf{F}(s,t;n,k)=s.$
By Lemma 2.10,
$\mathbf{F}(s_{1},t_{1};n-1,k)+\mathbf{F}(s,t_{2};k+1,k)-s_{1}\geq s_{1}+s-s_{1}=s.$
So we consider $s>0$ and t>(kd1)(nk)t>(k-d-1)(n-k). Denote
$\mathbf{F}:=\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}.$
We have both
$\mathbf{F}\geq\mathbf{F}(s,t_{2};k+1,k)+\mathbf{F}(s_{1},t_{1};n-1,k)-s_{1}.$ (67)
and
$\mathbf{F}\geq\mathbf{F}(s_{1},t_{1};n-1,k).$ (68)
Write the canonical expression
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$ (69)
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $m\in\{0,\dots,n-k-1\}$, $\tau\in(0,d+2]$. We have
$\mathbf{F}(s,t;n,k)=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
Next, we analyze $\mathbf{F}(s,t_{2};k+1,k)$.
Case (1): $t_{2}\leq k-d-1$We use
$\mathbf{F}\geq\mathbf{F}(s_{1},t_{1};n-1,k)\geq\mathbf{F}(s,t_{1};n-1,k).$
Since $t_{2}\leq k-d-1$, we have
$t_{1}\geq(k-d-1)(n-k-1)+(d+2)m+\tau.$
Therefore,
$\mathbf{F}\geq\mathbf{F}(s,(k-d-1)(n-k-1)+(d+2)m+\tau;n-1,k)=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
Therefore,
$\mathbf{F}\geq\mathbf{F}(s,t;n,k).$
Case (2): $t_{2}>k-d-1$We write $t_{2}=k-d-1+\tau_{2}$, where $\tau_{2}\in(0,d+2]$. We have
$\mathbf{F}(s,t_{2};k+1,k)=s+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}.$
Since $s_{1}\geq s$, write
$s_{1}=d+\sigma_{1},$
with $\sigma_{1}\geq\sigma$.
Subcase (2.1): $\sigma_{1}\in[\sigma,1]$We have
$t_{1}=t-t_{2}=(k-d-1)(n-k-1)+(d+2)m+\tau-\tau_{2}.$
Subsubcase (2.1.1): $\tau-\tau_{2}>0$We have the canonical expression for $(s_{1},t_{1})$:
$\begin{cases}s_{1}=d+\sigma_{1},\\ t_{1}=t-t_{2}=(k-d-1)(n-k-1)+(d+2)m+\tau-\tau_{2}.\end{cases}$
Therefore,
$\mathbf{F}(s_{1},t_{1};n-1,k)=s_{1}+m+\min\{\tau-\tau_{2},\frac{1}{2}(\sigma_{1}+\tau-\tau_{2}),1\}.$
Therefore,
$\mathbf{F}\geq s+m+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{\tau-\tau_{2},\frac{1}{2}(\sigma+\tau-\tau_{2}),1\}.$
We just need to show
$\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{\tau-\tau_{2},\frac{1}{2}(\sigma+\tau-\tau_{2}),1\}\geq\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
To show $\min_{i}a_{i}+\min_{j}b_{j}\geq\min_{k}c_{k}$, it suffices to show $\forall i,j,\exists k$, s.t. $a_{i}+b_{j}\geq c_{k}$. We can check:
$\bullet$ $\tau_{2}+\tau-\tau_{2}\geq\tau$.
$\bullet$ $\tau_{2}+\frac{1}{2}(\sigma+\tau-\tau_{2})\geq\frac{1}{2}(\sigma+\tau)$.
$\bullet$ $\frac{1}{2}(\sigma+\tau_{2})+\tau-\tau_{2}\geq\frac{1}{2}(\sigma+\tau).$
$\bullet$ $\frac{1}{2}(\sigma+\tau_{2})+\frac{1}{2}(\sigma+\tau-\tau_{2})\geq\frac{1}{2}(\sigma+\tau)$.
Subsubcase (2.1.2): $\tau-\tau_{2}\leq 0,m\geq 1$We write $t_{1}=(k-d-1)(n-k-1)+(d+2)(m-1)+d+2+\tau-\tau_{2}$. We have
$\mathbf{F}(s_{1},t_{1};n-1,k)=s_{1}+m-1+\min\{\tau-\tau_{2}+d+2,\frac{1}{2}(\sigma_{1}+\tau-\tau_{2}+d+2),1\}.$
Therefore,
$\mathbf{F}\geq s+m-1+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{\tau-\tau_{2}+d+2,\frac{1}{2}(\sigma_{1}+\tau-\tau_{2}+d+2),1\}.$
We just need to prove
$\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{\tau-\tau_{2}+d+1,\frac{1}{2}(\sigma+\tau-\tau_{2}+d),0\}\geq\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
This is not hard to verify by noting $\tau\leq\tau_{2}\leq d+2$.
$\bullet$ $\tau_{2}+\tau-\tau_{2}+d+1\geq\tau$.
$\bullet$ $\tau_{2}+\frac{1}{2}(\sigma+\tau-\tau_{2}+d)\geq\frac{1}{2}(\sigma+\tau)$.
$\bullet$ $\tau_{2}+0\geq\tau$.
$\bullet$ $\frac{1}{2}(\sigma+\tau_{2})+\tau-\tau_{2}+d+1\geq\tau+d+1-\frac{1}{2}\tau_{2}\geq\tau$.
$\bullet$ $\frac{1}{2}(\sigma+\tau_{2})+\frac{1}{2}(\sigma+\tau-\tau_{2}+d)\geq\frac{1}{2}(\sigma+\tau)$.
$\bullet$ $\frac{1}{2}(\sigma+\tau_{2})+0\geq\frac{1}{2}(\sigma+\tau)$.
$\bullet$ $1+\tau-\tau_{2}+d+1\geq\tau$.
$\bullet$ $1+\frac{1}{2}(\sigma+\tau-\tau_{2}+d)\geq\frac{1}{2}(\sigma+\tau)$.
$\bullet$ $1+0\geq 1$.
Subsubcase (2.1.3): $\tau-\tau_{2}\leq 0,m=0$We have $t_{1}\leq(k-d-1)(n-k-1)$, so
$\mathbf{F}(s_{1},t_{1};n-1,k)=s_{1}.$
Therefore,
$\mathbf{F}\geq\mathbf{F}(s,t_{2};k+1,k)=s+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}.$
Also, $\mathbf{F}(s,t;n,k)=s+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}$. We see that $\mathbf{F}\geq\mathbf{F}(s,t;n,k)$ since $\tau_{2}\geq\tau$.
Subcase(2.2): $\sigma_{1}>1$Write $s_{1}=(d+\Delta)+(\sigma_{1}-\Delta),$ where $\Delta\in\mathbb{N}^{+}$ so that $\sigma_{1}-\Delta\in(0,1]$. We have
$t_{1}$ $=(k-d-1)(n-k-1)+(d+2)m+\tau-\tau_{2}$
$=(k-d-\Delta-1)(n-k-1)+(d+2)m+\tau-\tau_{2}+\Delta(n-k-1)$
$=(k-d-\Delta-1)(n-k-1)+(d+2)m_{1}+\tau_{1}.$
Here, the last line is the canonical expression for $t_{1}$ (see (9)), in which $0\leq m_{1}\leq n-k-2$, $\tau_{1}\in(0,d+\Delta+2]$.
If $(d+2)m+\tau-\tau_{2}\leq 0$, then $m=0,\tau_{2}\geq\tau$. By the same reasoning as in Subsubcase (2.1.3), we have $\mathbf{F}\geq\mathbf{F}(s,t;n,k)$.
Hence, we can assume $(d+2)m+\tau-\tau_{2}>0$ We have
$\mathbf{F}(s_{1},t_{1};n-1,k)=s_{1}+m_{1}+\min\{\tau_{1},\frac{1}{2}(\sigma_{1}-\Delta+\tau_{1}),1\}.$
Since $(d+2)m+\tau-\tau_{2}\leq(d+2)m_{1}+\tau_{1}$ and $\tau,\tau_{1},\tau_{2}\in(0,d+2]$, we have $m_{1}\geq m-1$. If $m_{1}\geq m+1$, then $\mathbf{F}\geq\mathbf{F}(s_{1},t_{1};n-1,k)\geq s+m+1\geq\mathbf{F}(s,t;n,k)$. We only need to consider two cases: $m_{1}=m,m_{1}=m-1$.
We have
$\mathbf{F}\geq s+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+m_{1}+\min\{\tau_{1},\frac{1}{2}(\sigma_{1}-\Delta+\tau_{1}),1\}.$
Subsubcase (2.2.1): When $m_{1}=m,\tau_{1}=\tau-\tau_{2}+\Delta(n-k-1)$We have
$\mathbf{F}\geq s+m+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{(\tau-\tau_{2})\vee 0,\frac{1}{2}(\sigma+\tau-\tau_{2}),1\},$
where we used $\tau_{1}\geq(\tau-\tau_{2})\vee 0,$ and $\sigma_{1}-\Delta+\tau_{1}\geq\sigma-\Delta+\tau-\tau_{2}+\Delta(n-k-1)\geq\sigma+\tau-\tau_{2}$.
We just need to check
$\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{(\tau-\tau_{2})\vee 0,\frac{1}{2}(\sigma+\tau-\tau_{2}),1\}\geq\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
A less obvious case will be $\frac{1}{2}(\sigma+\tau_{2})+(\tau-\tau_{2})\vee 0\geq\frac{1}{2}(\sigma+\tau_{2})+\frac{1}{2}(\tau-\tau_{2})\geq\frac{1}{2}(\sigma+\tau)$. We leave out the verification of other cases.
Subsubcase (2.2.2): When $m_{1}=m-1,\tau_{1}=\tau-\tau_{2}+d+2+\Delta(n-k-1)$Since $n\geq k+2$, we have
$\mathbf{F}\geq s+m-1+\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{\tau-\tau_{2}+d+2+\Delta,\frac{1}{2}(\sigma+\tau-\tau_{2}+d+2),1\}.$
Since $\tau_{1}\leq d+2$, we have $\tau_{2}\geq\tau$. Note that
$\min\{\tau-\tau_{2}+d+2+\Delta,\frac{1}{2}(\sigma+\tau-\tau_{2}+d+2),1\}-1\geq\min\{(\tau-\tau_{2}+1)\vee 0,\frac{1}{2}(\sigma+\tau-\tau_{2}),0\}.$
We just need to prove
$\min\{\tau_{2},\frac{1}{2}(\sigma+\tau_{2}),1\}+\min\{(\tau-\tau_{2}+1)\vee 0,\frac{1}{2}(\sigma+\tau-\tau_{2}),0\}\geq\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.$
This is not hard to verify by noting $\tau_{2}\geq\tau$. A less obvious case is $\frac{1}{2}(\sigma+\tau_{2})+(\tau-\tau_{2}+1)\vee 0\geq\frac{1}{2}(\sigma+\tau_{2})+\frac{1}{2}(\tau-\tau_{2}+1)\geq\frac{1}{2}(\sigma+\tau)$.
Proof of Theorem 1.8.
Finally, we see that Theorem 1.8 can be obtained from Proposition 4.3 and Proposition 5.1 using induction on $(n,k)$. ∎
6 Exceptional set estimate: lower bound
We prove Theorem 1.10 in this section. The result in this section has been studied by the author [8] under the Euclidean setting, and the statement there is slightly different from this paper. We still provide the details for the finite-field setting, though they are essentially the same.
Proof of Theorem 1.10.
Recall that $(n,k)=(2,1)$ and $\frac{a}{2}<s\leq\min\{1,a\}$ was considered in Proposition 2.6. We denote $\mathbf{A}_{s,a}\subset\mathbb{F}_{p}^{2}$ to be the set constructed there which satisfies $\mathbf{A}_{s,a}\gtrsim p^{a}$, and
$E_{s}(\mathbf{A}_{s,a};2,1)\gtrsim p^{2s-a}.$
Next, we prove for general $(n,k)$. Note that Type 1 and Type 4 are easy, so we consider Type 2 and Type 3. Recall the canonical expression as in Definition 2.4: $a=m+\beta,s=l+\gamma$.
Type 2We have $l+1\leq m\leq n+l-k$. For $\gamma$, we only use $\gamma>\beta$. We first show the existence of $A$ with $A\sim p^{a}$ such that $E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m-l)(k-l)}$.
Choose $A=\mathbb{F}_{p}^{m}\times I\times\mathbf{0}^{n-m-1}$, where $I\subset\mathbb{F}_{p}$ has cardinality $p^{\beta}$. For simplicity, we just write $A=\mathbb{F}_{p}^{m}\times I$. We want to find those $V\in G(n-k,\mathbb{F}_{p}^{n})$ such that $(\pi_{V}^{*}(A))<p^{l+\gamma}$. We note that if $V$ satisfies $\pi_{V}^{*}(\mathbb{F}_{p}^{m})\leq p^{l}$, then $\pi_{V}^{*}(\mathbb{F}_{p}^{m}\times I)\leq p^{l+\beta}<p^{l+\gamma}$, which means $V\in E_{s}(A;n,k)$. This shows that
$E_{s}(A;n,k)\supset\{V\in G(n-k,n):\#\pi_{V}^{*}(\mathbb{F}_{p}^{m})\leq p^{l}\}.$
By Lemma 2.8, the right hand side has cardinality $\sim p^{k(n-k)-(k-l)(m-l)}$.
Type 3, $\gamma\in(\frac{\beta}{2},\beta]$We have $n-m-1\geq k-l,\gamma>\frac{\beta}{2}$. We show the existence of $A$ with $A\sim p^{a}$ such that $E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)+2\gamma-\beta}$.
We choose $A=\mathbf{A}_{\gamma,\beta}\times\mathbb{F}_{p}^{m}\times\mathbf{0}^{n-m-2}$, where $\mathbf{A}_{\gamma,\beta}\subset\mathbb{F}_{p}^{2}$ is such that $\mathbf{A}_{\gamma,\beta}\sim p^{\beta}$ and $E_{\gamma}(\mathbf{A}_{\gamma,\beta};2,1)\gtrsim p^{2\gamma-\beta}$. We will use $\mathbb{F}_{p}^{2}$ to denote $\mathbb{F}_{p}^{2}\times\mathbf{0}^{n-2}$. We use $\mathbb{F}_{p}^{m}$ to denote $\mathbf{0}^{2}\times\mathbb{F}_{p}^{m}\times\mathbf{0}^{n-m-2}$. We use $\theta$ to denote the lines in $G(1,\mathbb{F}_{p}^{2})$. Define
$\mathcal{U}_{\theta}=\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(\pi^{*}_{V}(\theta\oplus\mathbb{F}_{p}^{m}))\leq p^{l},\#(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))=p^{l},\#(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))=p^{l+1}\},$
$\mathcal{V}_{\theta}=\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))\leq p^{l}\},$
$\mathcal{W}_{\theta}=\{V\in\mathcal{V}_{\theta}:\#(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))\leq p^{l-1}\}\cup\{V\in\mathcal{V}_{\theta}:\#(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m})\leq p^{l}\}.$
It is not hard to see $\mathcal{U}_{\theta}=\mathcal{V}_{\theta}\setminus\mathcal{W}_{\theta}$.
Let $\mathbf{E}_{\gamma,\beta}:=E_{\gamma}(\mathbf{A}_{\gamma,\beta};2,1)=\{\theta\in G(1,\mathbb{F}_{p}^{2}):\#(\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta}))<p^{\gamma}\}$. We claim that if $\theta\in\mathbf{E}_{\gamma,\beta}$, then $\mathcal{V}_{\theta}\subset E_{l+\gamma}(A)$. In other words, $V\in\mathcal{V}_{\theta}$ satisfies $(\pi_{V}^{*}(A))<p^{l+\gamma}$. We first note that by definition $\mathbf{A}_{\gamma,\beta}\subset\bigcup_{L\in\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta})}L$, so
$A=\mathbf{A}_{\gamma,\beta}\times\mathbb{F}_{p}^{m}\subset(\bigcup_{L\in\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta})}L)\times\mathbb{F}_{p}^{m}=(\bigcup_{L\in\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta})}L\times\mathbb{F}_{p}^{m}).$
Therefore,
$(\pi_{V}^{*}(A))\leq\#(\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta}))\cdot\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))<p^{\gamma+l}.$
We see that $E_{l+\gamma}(A;n,k)\supset\bigcup_{\theta\in\mathbf{E}_{\gamma,\beta}}\mathcal{V}_{\theta}.$ $\mathcal{V}_{\theta}$ may intersect with each other, so we delete a tiny portion $\mathcal{W}_{\theta}$ from $\mathcal{V}_{\theta}$, so that the remaining part $\mathcal{U}_{\theta}=\mathcal{V}_{\theta}\setminus\mathcal{W}_{\theta}$ are disjoint with each other. Then we have
$E_{l+\gamma}(A;n,k)\supset\bigsqcup_{\theta\in\mathbf{E}_{\gamma,\beta}}\mathcal{U}_{\theta}.$
To make the argument precise, we first apply Lemma 2.8 to see
$\mathcal{V}_{\theta}\sim p^{k(n-k)-(m+1-l)(k-l)}.$
To calculate $\mathcal{W}_{\theta}$, we apply Lemma 2.8 again to see
$\{V\in G(n-k,n):\dim(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))\leq p^{l-1}\}\sim p^{k(n-k)-(m-l+1)(k-l+1)},$
and
$\{V\in G(n-k,n):\dim(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))\leq p^{l}\}\sim p^{k(n-k)-(m+2-l)(k-l)}.$
Therefore, $\mathcal{W}_{\theta}\ll\#\mathcal{V}_{\theta}$ (when $p$ is large enough) and hence
$\mathcal{U}_{\theta}\sim\#\mathcal{V}_{\theta}\sim p^{k(n-k)-(m+1-l)(k-l)}.$
Next we show $\mathcal{U}_{\theta}\cap\mathcal{U}_{\theta^{\prime}}=\emptyset$ for $\theta\neq\theta^{\prime}$. By contradiction, suppose $V\in\mathcal{U}_{\theta}\cap\mathcal{U}_{\theta^{\prime}}$. By definition, $V$ satisfies
$\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))\leq p^{l},$
$\#(\pi_{V}^{*}(\theta^{\prime}\oplus\mathbb{F}_{p}^{m}))\leq p^{l},$
$\#(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))=p^{l},$
$\#(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))=p^{l+1}.$
The first and third conditions imply $\pi_{V}^{*}(\theta)\subset\pi_{V}^{*}(\mathbb{F}_{p}^{m})$. Similarly, the second and third conditions imply $\pi_{V}^{*}(\theta^{\prime})\subset\pi_{V}^{*}(\mathbb{F}_{p}^{m})$. Noting that $\textup{span}\{\theta,\theta^{\prime}\}=\mathbb{F}_{p}^{2}$, we have
$\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m})\subset\pi_{V}^{*}(\mathbb{F}_{p}^{m}),$
which has cardinality $p^{l}$, contradicting the fourth condition.
We have shown that $E_{l+\gamma}(A;n,k)\supset\bigsqcup_{\theta\in\mathbf{E}_{\gamma,\beta}}\mathcal{U}_{\theta}$, which has cardinality $\mathbf{E}_{\gamma,\beta}\cdot\#\mathcal{U}_{\theta}$
$\gtrsim p^{2\gamma-\beta+k(n-k)-(m+1-l)(k-l)}.$
Type 2, $\gamma\in(\frac{\beta+1}{2},1]$The trick is to write $a=(m-1)+(\beta+1)=:m^{\prime}+\beta^{\prime}$. Note that $l\leq m^{\prime}\leq n+l-k-1$ and $\gamma>\frac{\beta^{\prime}}{2}$. We apply the construction in “Type 3, $\gamma\in(\frac{\beta}{2},\beta]$”, where we only used $\gamma>\frac{\beta}{2}$. We obtain a $A$ with $A\sim p^{a}$ such that
$E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m^{\prime}+1-l)(k-l)+2\gamma-\beta^{\prime}}=p^{k(n-k)-(m-l)(k-l)+2\gamma-(\beta+1)}.$
Type 3, $\gamma\in(0,\frac{\beta}{2}]$The trick is to make $a$ bigger. We write $a=m+\beta\leq(m+1)+0=:m^{\prime}+\beta^{\prime}$. One can check $l+1\leq m^{\prime}\leq n+l-k$ and $\gamma>\beta^{\prime}$. We apply the construction in Type 2 with $(m^{\prime},\beta^{\prime})$. We find $A^{\prime}$ with $A^{\prime}\sim p^{m+1}$ such that
$E_{s}(A^{\prime};n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)}.$
Choosing any subset $A\subset A^{\prime}$ with $A\sim p^{a}$, we get
$E_{s}(A;n,k)\geq\#E_{s}(A^{\prime};n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)}.$
7 Exceptional set estimate: Upper bound I
We prove the following proposition in this section.
Proposition 7.1.
Assume Conjecture 1.5 holds. For $0<a\leq k+1$ and $0<s$, and $A\subset\mathbb{F}_{p}^{k+1}$ with $A\gtrsim p^{a}$, we have
$E_{s-\varepsilon}(A;k+1,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;k+1,k)},$ (70)
for any $\varepsilon>0$.
It is proved by using the following lemma.
Lemma 7.2.
Assume Conjecture 1.5 holds. Fix $0<\varepsilon_{0}<1/100$. Let $0<a\leq k+1,s>0$ and $\sigma\geq 0$. Suppose that the parameters satisfy $\mathbf{M}(a,s;k+1,k)+\varepsilon_{0}\leq k$ and $a-s+\varepsilon_{0}\leq\sigma\leq a$. Let $A\subset\mathbb{F}_{p}^{k+1}$ with $A=p^{k+1}$. Define
$\mathcal{V}:=\{L\in A(1,\mathbb{F}_{p}^{k+1}):p^{\sigma}\leq\#(L\cap A)\}.$
Then
$\mathcal{V}\lesssim_{\varepsilon_{0}}p^{\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}}.$
Remark 7.3.
The implicit constant in $\lesssim_{\varepsilon_{0}}$ can be made independent of $\sigma$, since we can apply the lemma to $\sigma\in\varepsilon_{0}^{2}\cdot\mathbb{N}$ which consists of $O(\varepsilon_{0}^{-2})$ values.
Proof.
The proof of this part is by a Furstenberg estimate. As in Definition 2.4, write
$\begin{cases}a=m+\beta\\ s=l+\gamma.\end{cases}$ (71)
There are four cases.
Type 1: When $s>\min\{a,k\}$,
$\mathbf{M}(a,s;k+1,k)=k.$
Type 2: When $s\leq\min\{a,k\}$ and $m=l+1,\gamma\in(\beta,1]$,
$\mathbf{M}(a,s;k+1,k)=l+\max\{2\gamma-(\beta+1),0\}.$
Type 3: When $s\leq\min\{a,k\}$ and $m=l,\gamma\in(0,\beta]$,
$\mathbf{M}(a,s;k+1,k)=l+\max\{2\gamma-\beta,0\}.$
Type 4: When $s\leq a-1$,
$\mathbf{M}(a,s;k+1,k)=-\infty.$
Type 1 is trivial, since $\mathcal{V}\leq p^{k}$ is always true. Type 4 is also easy, since in this case we have $\sigma\geq a-s+\varepsilon_{0}\geq 1+\varepsilon_{0}$, and hence $\mathcal{V}=\emptyset$. We only need to consider Type 2 and Type 3.
Suppose $\mathcal{V}=p^{t}$, we want to show
$t\leq\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}.$
We write $t=2w+\tau$, where $w\in\{0,\dots,k-1\}$, $\tau\in(0,2]$.
For convenience, we define
$R(\beta,\gamma):=\begin{cases}\max\{2\gamma-(\beta+1),0\}&m=l+1,\gamma\in(\beta,1]\\ \max\{2\gamma-\beta,0\}&m=l,\gamma\in(0,\beta].\end{cases}$ (72)
Then $\mathbf{M}(a,s;k+1,k)=l+R(\beta,\gamma)$ in the case of Type 2 or Type 3.
Note that $A$ contains a $(\sigma,t;k+1,1)$-Furstenberg set $\bigcup_{L\in\mathcal{V}}(L\cap A)$. Since Conjecture 1.5 holds, by Theorem 1.8, we have the lower bound:
$a+\eta_{0}\geq\mathbf{F}(\sigma,t;k+1,1)=w+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\},$ (73)
for any $\eta_{0}>0$ when $p$ is big enough depending on $\eta_{0}$. We choose $\eta_{0}$ such that $\eta_{0}<\varepsilon_{0}/2$
From (73), there are three possible cases.
(1) $a+\eta_{0}\geq w+\sigma+\tau$.
This implies $t=2w+\tau\leq a+w-\sigma+\eta_{0}$. Our goal is to show $t\leq\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}$, so we just need to show
$\mathbf{M}(a,s;k+1,k)\geq w.$
We also know
$a\geq w+\sigma+\tau-\eta_{0}\geq w+\tau+a-s+\varepsilon_{0}-\eta_{0}\geq w+\tau+a-s+\varepsilon_{0}/2,$
which implies
$s\geq w+\tau+\varepsilon_{0}/2.$
Since $s\leq l+1$ and $w\in\mathbb{N}$, we have $w\leq l$. Therefore, with the condition $\mathbf{M}(a,s;k+1,k)\geq l$, we see that
$\mathbf{M}(a,s;k+1,k)\geq w.$
(2) $a+\eta_{0}\geq\sigma+1+w$.
In this case, $a\geq\sigma+1+w-\eta_{0}\geq a-s+w+1+\varepsilon_{0}-\eta_{0}$, which implies
$s\geq w+1+\varepsilon_{0}/2.$
Since $s\leq l+1$ and $w\in\mathbb{N}$, we get $w\leq l-1$.
Now we want to prove $t\leq\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}$. If this is not true, then by contradiction, we assume
$2w+\tau=t>\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}.$
This will imply
$l-1+\tau\geq w+\tau>\mathbf{M}(a,s;k+1,k)+a-\sigma-w+\varepsilon_{0}\geq\mathbf{M}(a,s;k+1,k)+1+\varepsilon_{0}-\eta_{0}\geq l+1,$
which is a contradiction, as $\tau\leq 2$.
(3) $a+\eta_{0}\geq w+\frac{3}{2}\sigma+\frac{1}{2}\tau$.
This is equivalent to $t\leq 2a-3\sigma+2\eta_{0}$. We hope to prove
$t\leq\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}.$
We only need to prove
$2a-3\sigma\leq\mathbf{M}(a,s;k+1,k)+a-\sigma,$
since $2\eta_{0}<\varepsilon_{0}$.
This is equivalent to
$a\leq\mathbf{M}(a,s;k+1,k)+2\sigma.$
We plug in $a=m+\beta,s=l+\gamma,\mathbf{M}(a,s;k+1,k)=l+R(\beta,\gamma),\sigma\geq(m+\beta)-(l+\gamma)+\varepsilon_{0}$. Then it suffices to prove
$m+\beta\leq l+R(\beta,\gamma)+2(m+\beta)-2(l+\gamma)+2\varepsilon_{0}.$
This inequality is equivalent to
$2\gamma-(\beta+m-l)\leq R(\beta,\gamma)+2\varepsilon_{0},$
This is easy to verify by checking either in the case $m=l+1,\gamma\in(\beta,1]$ or in the case $m=l,\gamma\in(0,\beta]$. ∎
Proof of Proposition 7.1.
We will use the lemma that we just proved. We show that if $A\subset\mathbb{F}_{p}^{k+1}$ with $A\gtrsim p^{a}$, $E\subset G(1,\mathbb{F}_{p}^{k+1})$ with $E>C_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;k+1,k)}$, and $p$ is big enough, then there exists $V\in E$ such that
$\pi_{V}^{*}(A)>p^{s-\varepsilon}.$
Since $E\leq p^{k}$, we have
$\varepsilon+\mathbf{M}(a,s;k+1,k)\leq k.$
For each $V\in E$ and a dyadic number $1\leq\mu\leq p$, define
$\mathbb{L}_{V,\mu}:=\{L\in A(1,\mathbb{F}_{p}^{k+1}):L\parallel V,\mu\leq\#(A\cap L)<2\mu\}.$
We see that $\{\mathbb{L}_{V,\mu}\}_{\mu}$ form a partition of the lines parallel to $V$.
We choose $\varepsilon_{0}\ll\varepsilon$, for example, $\varepsilon_{0}=\varepsilon^{2}$. We will estimate the cardinality of the set $\bigcup_{V\in E}\mathbb{L}_{V,\mu}$ for $\mu\geq p^{a-s+\varepsilon_{0}}$. We write $\mu=p^{\sigma}$ with $a\geq\sigma\geq a-s+\varepsilon_{0}$. By Lemma 7.2, we have
$\bigcup_{V\in E}\mathbb{L}_{V,\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+a-\sigma+\varepsilon_{0}}.$
If we define
$E_{\mu}:=\{V\in E:\#\mathbb{L}_{V,\mu}>p^{a-\sigma-\varepsilon_{0}}\},$
then
$E_{\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+2\varepsilon_{0}}.$
Summing over dyadic $\mu\geq p^{a-s+\varepsilon_{0}}$, we have
$\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}\lesssim_{\varepsilon}p^{\mathbf{M}(a,s;k+1,k)+3\varepsilon_{0}}.$
Therefore, $E>\#\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}$ when $p$ is large enough. We can find $V\in E\setminus\bigcup_{\mu\geq p^{a-s+\varepsilon_{0}}}E_{\mu}$. In other words, there exists $V\in E$, such that for any $\mu=p^{\sigma}\geq p^{a-s+\varepsilon_{0}}$,
$\mathbb{L}_{V,\mu}\leq p^{a-\sigma-\varepsilon_{0}}.$
Now, we do the estimate. We have
$A=\sum_{\mu<p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L)+\sum_{\mu\geq p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L).$
Note that
$\sum_{\mu\geq p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L)\leq\sum_{\mu\geq p^{a-s+\varepsilon_{0}}}2\mu\cdot\#\mathbb{L}_{V,\mu}\leq 2\sum_{\mu\geq p^{a-s+\varepsilon_{0}}}p^{a-\varepsilon_{0}}\leq\frac{1}{2}\#A.$
In the last inequality, we are summing over $O(\log p)$ terms and we assume $p$ is big enough.
We have
$\frac{1}{2}\#A\leq\sum_{\mu<p^{a-s+\varepsilon_{0}}}\sum_{L\in\mathbb{L}_{V,\mu}}\#(A\cap L)<2p^{a-s+\varepsilon_{0}}\#\pi_{V}^{*}(A).$
This implies $\pi_{V}^{*}(A)\gtrsim p^{s-\varepsilon_{0}}>p^{s-\varepsilon}$.
8 Exceptional set estimate: Upper bound II
The goal of this section is to prove the following proposition.
Proposition 8.1.
Let $n,k\in\mathbb{N}$ and $n\geq k+2$. Suppose for any $n^{\prime}=k+1,\dots,n-1$ and $a\in(0,n^{\prime}],s>0$, we have
$\sup_{A\subset\mathbb{F}_{p}^{n^{\prime}},\#A\sim p^{a}}\#E_{s-\varepsilon}(A;n^{\prime},k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n^{\prime},k)},$ (74)
for any $\varepsilon>0$. Then for $a\in(0,n],s>0$, we have
$\sup_{A\subset\mathbb{F}_{p}^{n},\#A\sim p^{a}}\#E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n,k)},$
for any $\varepsilon>0$.
We prove the following lemma.
Lemma 8.2.
Assume (74) holds. Suppose $n,k\in\mathbb{N}$, $n\geq k+2$, $a\in(0,n]$, $s>0$ and $s\in(a-(n-k),\min\{a,k\}]$. Then for any $A\subset\mathbb{F}_{p}^{n}$ with $A\sim p^{a}$, there exist numbers $a_{1},s_{1}$, such that
$\begin{cases}p^{\max\{0,a-1\}}\leq p^{a_{1}}\leq p^{\min\{n-1,a\}};\\ 1\leq p^{s_{1}}\leq p^{s};\end{cases}$ (75)
and
$E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)}.$ (76)
Proof.
We will do several steps of pigeonholing, which lead to the parameters $s_{1},a_{1}$. For convenience, we denote $\mathcal{V}:=E_{s-\varepsilon}(A;n,k)$.
By properly choosing the coordinate, we can assume $\gtrsim 1$ fraction of the (nk)(n-k)-planes in $\mathcal{V}$ intersect $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$ transversely (at a (nk1)(n-k-1)-plane). Since we are going to find upper bound for $\mathcal{V}$, we can just assume all the planes in $\mathcal{V}$ intersect $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$ at a (nk1)(n-k-1)-plane.
For each $W\in G(n-k-1,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$, we define $\mathcal{V}_{W}:=\{V\in\mathcal{V}:V\cap(\mathbb{F}_{p}^{n-1}\times\mathbf{0})=W\}$. By pigeonholing, there exits $\mathcal{W}\subset G(n-k-1,\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with $\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, $\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $p^{t_{1}},p^{t_{2}}\geq 1$ are dyadic numbers so that
$p^{t_{1}+t_{2}}\gtrapprox\#\mathcal{V}.$
By pigeonholing on $A\cap(\mathbb{F}_{p}^{n-1}\times\xi)$ for $\xi\in\mathbb{F}_{p}$, we can find a dyadic number $1\leq p^{a_{1}}\leq p^{a}$ and $\Xi_{1}\subset\mathbb{F}_{p}$ so that $A\cap(\mathbb{F}_{p}^{n-1}\times\xi)\sim p^{a_{1}}$ for any $\xi\in\Xi_{1}$ and
$p^{a_{1}}\cdot\#\Xi_{1}\gtrapprox p^{a}.$
Since $\Xi_{1}\leq p$, we also have $p^{a_{1}}\gtrapprox p^{a-1}$.
For each $\xi\in\Xi_{1}$, we consider the slice $\mathbb{F}_{p}^{n-1}\times\xi$. We denote $A_{\xi}:=A\cap(\mathbb{F}_{p}^{n-1}\times\xi)$. There exists a dyadic number $1\leq p^{s_{1}(\xi)}\leq p^{s-\varepsilon}$, such that
$\pi_{W}^{*}(A_{\xi})\sim p^{s_{1}(\xi)},$
for $\gtrapprox\#\mathcal{W}\sim p^{t_{1}}$ many $W\in\mathcal{W}$. Here, $s_{1}(\xi)\leq s-\varepsilon$ is because $\pi_{W}^{*}(A_{\xi})\leq\#\pi_{V}^{*}(A)<p^{s-\varepsilon}$ for $V\in\mathcal{V}_{W}$.
We see that $E_{s_{1}(\xi)+\varepsilon^{2}/2}(A_{\xi};n-1,k)$ contains a $\gtrapprox 1$ portion of $\mathcal{W}$ when $p$ is large enough depending on $\varepsilon$. We can use (74) for the quadruple $(a_{1},s_{1}(\xi)+\varepsilon^{2};n-1,k)$ to obtain
$p^{t_{1}}\lessapprox\#E_{s_{1}(\xi)+\varepsilon^{2}/2}(A_{\xi};n-1,k)\lesssim_{\varepsilon}p^{\varepsilon^{2}/2+\mathbf{M}(a_{1},s_{1}(\xi)+\varepsilon^{2};n-1,k)}.$
By pigeonholing on $s_{1}(\xi)$, we can find $\Xi\subset\Xi_{1}$ such that
$s_{1}(\xi)=s_{1},$
for all $\xi\in\Xi$. And,
$\Xi\gtrapprox\#\Xi_{1}\gtrapprox p^{a-a_{1}}.$
Next, we note that
$\sum_{W\in\mathcal{W}}\sum_{\xi\in\Xi}\#\pi_{W}^{*}(A_{\xi})\gtrapprox\#\mathcal{W}\cdot\#\Xi\cdot p^{s_{1}}.$
By pigeonholing, there exists $W\in\mathcal{W}$ such that
$\sum_{\xi\in\Xi}\#\pi_{W}^{*}(A_{\xi})\gtrapprox\#\Xi\cdot p^{s_{1}}\gtrapprox p^{a-a_{1}+s_{1}}.$
We fix this $W$ and study the behavior of $\mathcal{V}_{W}$. Without loss of generality, we can assume $W=\mathbb{F}_{p}^{n-1-k}\times\mathbf{0}^{k}\times\mathbf{0}$. For $V\in\mathcal{V}_{W}$, then $V\supset W$. We see that $V$ is of form $V=\mathbb{F}_{p}^{n-1-k}\times\ell_{V}$, where $\ell_{V}\in G(1,\mathbf{0}^{n-1-k}\times\mathbb{F}_{p}^{k+1})$. Since $V\cap\mathbb{F}_{p}^{n-1}\times\mathbf{0}=W$, we have that $\ell_{V}$ is not parallel to $\mathbb{F}_{p}^{n-1}\times\mathbf{0}$.
Next, we project everything onto $\mathbf{0}^{n-1-k}\times\mathbb{F}_{p}^{k+1}$. For $x=(x_{1},\dots,x_{n})\in\mathbb{F}_{p}^{n}$, we define
$\pi_{\mathbb{F}_{p}^{k+1}}(x):=(0,\dots,0,x_{n-k},x_{n-k+1},\dots,x_{n}).$
One can also check that $\pi_{W}^{*}(x)=\pi_{\mathbb{F}_{p}^{k+1}}(x)+W$.
Define $B_{\xi}=\pi_{\mathbb{F}_{p}^{k+1}}(A_{\xi})$. We see that $B_{\xi}$ is a subset of $\mathbf{0}^{n-1-k}\times\mathbb{F}_{p}^{k+1}$ and $B_{\xi}=\#\pi_{W}^{*}(A_{\xi})$.
The next important observation is that for any $A\subset\mathbb{F}_{p}^{n}$, $V=\mathbb{F}_{p}^{n-1-k}\times\ell_{V}\in\mathcal{V}_{W}$, one has
$\pi_{V}^{*}(A)=\#\pi_{\ell_{V}}^{*}(\pi_{\mathbb{F}_{p}^{k+1}}(A)).$
This is not hard to verify once writing down the definition.
We can now do the estimate. By definition, $\pi_{V}^{*}(A)<p^{s-\varepsilon}$ for $V\in\mathcal{V}_{W}$. Therefore,
$p^{s-\varepsilon}>\#\pi_{\ell_{V}}^{*}(\pi_{\mathbb{F}_{p}^{k+1}}(A)).$
This means that
$\{\ell_{V}:V\in\mathcal{V}_{W}\}\subset E_{s-\varepsilon}(\pi_{\mathbb{F}_{p}^{k+1}}(A);k+1,k).$
Also, note that
$\pi_{\mathbb{F}_{p}^{k+1}}(A)\geq\#\pi_{\mathbb{F}_{p}^{k+1}}(\bigcup_{\xi\in\Xi}A_{\xi})=\#\pi^{*}_{W}(\bigcup_{\xi\in\Xi}A_{\xi})=\sum_{\xi\in\Xi}\#\pi_{W}^{*}(A_{\xi})\gtrapprox p^{a-a_{1}+s_{1}}.$
By (74) with $n^{\prime}=k+1$, we get
$p^{t_{2}}\lessapprox\#\mathcal{V}_{W}\leq\#E_{s-\varepsilon}(\pi_{\mathbb{F}_{p}^{k+1}}(A);k+1,k)\lesssim_{\varepsilon}p^{\varepsilon/3+\mathbf{M}(a-a_{1}+s_{1}-\varepsilon^{2},s-\varepsilon/2;k+1,k)}.$
(It is crucial that we use $s-\varepsilon/2$ instead of $s$ on the right hand side.)
Combining with the upper bound of $p^{t_{1}}$, we get
$\mathcal{V}\lessapprox p^{t_{1}+t_{2}}\lesssim_{\varepsilon}p^{\varepsilon/2+\mathbf{M}(a_{1},s_{1}+\varepsilon^{2};n-1,k)+\mathbf{M}(a-a_{1}+s_{1}-\varepsilon^{2},s-\varepsilon/2;k+1,k)}.$
Let $s_{1}^{\prime}=s_{1}+2\varepsilon^{2}$. Since $s_{1}\leq s-\varepsilon$, we have $s_{1}^{\prime}\leq s$.
Recall $3\varepsilon^{2}<\varepsilon/2$ and note $\mathbf{M}(a,s;n,k)$ is decreasing in $a$. By Lemma 2.15, we have
$\mathbf{M}(a-a_{1}+s_{1}^{\prime}-3\varepsilon^{2},s-\varepsilon/2;k+1,k)\leq\mathbf{M}(a-a_{1}+s_{1}^{\prime},s;k+1,k).$
Therefore, under the new $s_{1}^{\prime}$, we have
$\mathcal{V}\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1}^{\prime}-\varepsilon^{2};n-1,k)+\mathbf{M}(a-a_{1}+s_{1}^{\prime},s;k+1,k)}.$
Next, we want to slightly increase $a_{1}$ to $a_{1}^{\prime}$ so that $p^{\min\{0,a-1\}}\leq p^{a_{1}^{\prime}}$. By the condition $p^{\min\{0,a-1\}}\lessapprox p^{a_{1}}$, we have $p^{\min\{0,a-1\}}\leq p^{a_{1}+\varepsilon^{2}}$ when $p$ is large enough. We just need to let $a_{1}^{\prime}=\min\{a_{1}+\varepsilon^{2},\min\{0,a-1\}\}$. Noting $a_{1}^{\prime}\leq a_{1}+\varepsilon^{2}$ and by Lemma 2.15 again, we have
$\mathbf{M}(a_{1},s_{1}^{\prime}-\varepsilon^{2};n-1,k)\leq\mathbf{M}(a_{1}^{\prime}-\varepsilon^{2},s_{1}^{\prime}-\varepsilon^{2};n-1,k)\leq\mathbf{M}(a_{1}^{\prime},s_{1}^{\prime};n-1,k)$
We still use $s_{1},a_{1}$ for $s_{1}^{\prime},a_{1}^{\prime}$. We obtain
$\mathcal{V}\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(a-a_{1}+s_{1},s;k+1,k)}.$
Lemma 8.3.
Suppose $n,k\in\mathbb{N}$, $n\geq k+2$, $a\in(0,n]$, $s>0$ and $s\in(a-(n-k),\min\{a,k\}]$. Suppose there exist numbers $a_{1},s_{1}$, such that
$\begin{cases}\max\{0,a-1\}\leq a_{1}\leq\min\{n-1,a\};\\ 0\leq s_{1}\leq s.\end{cases}$ (77)
Then
$\mathbf{M}(a_{1},s_{1};n-1,k)+\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)\leq\mathbf{M}(a,s;n,k).$ (78)
Proof.
Write $a=m+\beta,s=l+\gamma$; $a_{1}=m_{1}+\beta_{1},s_{1}=l_{1}+\gamma_{1}$ in the canonical expressions. Since $a-1\leq a_{1}\leq a$, there are only two choices for $m_{1}$: $m_{1}=m$ or $m_{1}=m-1$. Since $s_{1}\leq s$, we have $l_{1}\leq l$.
Case 1: $m_{1}=m,l_{1}\leq l-1$
By Lemma 2.16,
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m-l+1)(k-l+1)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}.$
Trivially,
$\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)\leq k.$
Also, by Lemma 2.16,
$\mathbf{M}(a,s;n,k)\geq k(n-k)-(m+1-l)(k-l).$
It suffices to prove
$\max\{2\gamma_{1}-\beta_{1}-1,0\}\leq m+1-l.$
We just need to note $m\geq l$ and $2\gamma_{1}\leq 2$.
Case 2: $m_{1}=m,l_{1}=l$
Case 2.1: $\gamma>\beta,\gamma_{1}>\beta_{1}$
Since $s\leq a$ and $\gamma>\beta$, we have $m\geq l+1$. Hence,
$\begin{split}\mathbf{M}(a_{1},s_{1};n-1,k)&=\mathbf{M}(m+\beta_{1},l+\gamma_{1};n-1,k)\\ &=k(n-1-k)-(m-l)(k-l)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}.\end{split}$
Next, we estimate $\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)=\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)$. We claim that we have
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma_{1}+\beta-\beta_{1}\geq\gamma;\\ k&\gamma_{1}+\beta-\beta_{1}<\gamma.\end{cases}$ (79)
We just consider when $\gamma_{1}+\beta-\beta_{1}\geq\gamma$. We note $\gamma_{1}+\beta-\beta_{1}\in(0,2)$. If $\gamma_{1}+\beta-\beta_{1}\in(0,1]$, then we use Type 3 in Definition 2.4, which gives the desired formula. If $\gamma_{1}+\beta-\beta_{1}\in(1,2]$, we use Type 2, which gives $\mathbf{M}(l+1+(\gamma_{1}+\beta-\beta_{1}-1),l+\gamma;k+1,k)=k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}$, which is the same.
We also have $\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}$.
When $\gamma_{1}+\beta-\beta_{1}<\gamma$, we just need to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}\leq\max\{2\gamma-(\beta+1),0\},$
which is easy to verify, since $\gamma_{1}\leq\gamma$ and $\gamma_{1}+\beta-\beta_{1}<\gamma$.
When $\gamma_{1}+\beta-\beta_{1}\geq\gamma$, we just need to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq 1+\max\{2\gamma-(\beta+1),0\}.$
Just need to check three scenarios:
(1) $2\gamma_{1}-(\beta_{1}+1)\leq 1$, since $\gamma_{1}\leq 1$.
(2) $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 1+2\gamma-(\beta+1)$, since $\gamma_{1}>\beta_{1}$.
(3) Their sum $2\gamma_{1}-(\beta_{1}+1)+2\gamma-(\gamma_{1}+\beta-\beta_{1})=2\gamma+\gamma_{1}-(\beta+1)\leq 1+2\gamma-(\beta+1)$.
Case 2.2: $\gamma>\beta,\gamma_{1}\leq\beta_{1}$
In this case, by Lemma 2.16,
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m+1-l)(k-l)+\max\{2\gamma_{1}-\beta_{1},0\}.$
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k.$
$\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}.$
It suffices to prove
$\max\{2\gamma_{1}-\beta_{1},0\}\leq 1+\max\{2\gamma-(\beta+1),0\}.$
We just use $2\gamma_{1}-\beta_{1}\leq 1$, since $\gamma_{1}\leq\beta_{1}$.
Case 2.3: $\gamma\leq\beta,\gamma_{1}>\beta_{1}$
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m-l)(k-l)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}.$
(We use $\leq$ instead of ==, since when $m=l$, $\mathbf{M}(a_{1},s_{1};n-1,k)=k(n-1-k)$.)
Since $\gamma_{1}+\beta-\beta_{1}>\gamma$,
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}.$
Also, we have
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
It suffices to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq\max\{2\gamma-\beta,0\}.$
(1) $2\gamma_{1}-(\beta_{1}+1)\leq 2\gamma-\beta$, since $s_{1}\leq s$ implies $\gamma_{1}\leq\gamma$, and $\beta\leq 1$.
(2) $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 2\gamma-\beta$, since $\gamma_{1}>\beta_{1}$.
(3) $2\gamma_{1}-(\beta_{1}+1)+2\gamma-(\gamma_{1}+\beta-\beta_{1})=2\gamma+\gamma_{1}-(\beta+1)\leq 2\gamma-\beta$, since $\gamma_{1}\leq 1$.
Case 2.4: $\gamma\leq\beta,\gamma_{1}\leq\beta_{1}$
$\mathbf{M}(a_{1},s_{1};n-1,k)=k(n-1-k)-(m+1-l)(k-l)+\max\{2\gamma_{1}-\beta_{1},0\}.$
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)=\begin{cases}k&\gamma>\gamma_{1}+\beta-\beta_{1}\\ k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma\leq\gamma_{1}+\beta-\beta_{1}.\end{cases}$
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
When $\gamma>\gamma_{1}+\beta-\beta_{1}$, it suffices to show
$\max\{2\gamma_{1}-\beta_{1},0\}\leq\max\{2\gamma-\beta,0\}.$
This is true by noting $\gamma_{1}\leq\gamma$ and $\gamma>\gamma_{1}+\beta-\beta_{1}$.
When $\gamma\leq\gamma_{1}+\beta-\beta_{1}$, it suffices to show
$\max\{2\gamma_{1}-\beta_{1},0\}+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq 1+\max\{2\gamma-\beta,0\}.$
(1) $2\gamma_{1}-\beta_{1}\leq 1+2\gamma-\beta$, since $\gamma_{1}\leq\gamma$ and $\beta\leq 1$.
(2) $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 1+2\gamma-\beta$, since $\beta_{1}\leq 1$.
(3) $2\gamma_{1}-\beta_{1}+2\gamma-(\gamma_{1}+\beta-\beta_{1})=\gamma_{1}-\beta+2\gamma\leq 1+2\gamma-\beta$, since $\gamma_{1}\leq 1$.
Case 3: $m_{1}=m-1,l_{1}\leq l-2$The proof is similar to Case 1. By Lemma 2.16,
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m-l)(k-l+2)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}.$
Trivially,
$\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)\leq k.$
Also, by Lemma 2.16,
$\mathbf{M}(a,s;n,k)\geq k(n-k)-(m-l)(k-l+1).$
It suffices to prove
$\max\{2\gamma_{1}-\beta_{1}-1,0\}\leq m-l.$
We just need to note $m-l\geq 1$ and $2\gamma_{1}\leq 2$.
Case 4: $m_{1}=m-1,l_{1}=l-1$
Case 4.1: $\gamma>\beta,\gamma_{1}>\beta_{1}$
Since $s\leq a$ and $\gamma>\beta$, we have $m\geq l+1$. Hence,
$\begin{split}\mathbf{M}(a_{1},s_{1};n-1,k)&=\mathbf{M}(m-1+\beta_{1},l-1+\gamma_{1};n-1,k)\\ &=k(n-1-k)-(m-l)(k-l+1)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}.\end{split}$
Next, we estimate $\mathbf{M}(s_{1}+a-a_{1},s;k+1,k)=\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)$. As is Case 2.1,
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma_{1}+\beta-\beta_{1}\geq\gamma;\\ k&\gamma_{1}+\beta-\beta_{1}<\gamma.\end{cases}$ (80)
We also have $\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}$.
When $\gamma_{1}+\beta-\beta_{1}<\gamma$, we only need to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}\leq 1+\max\{2\gamma-(\beta+1),0\},$
which is easy to verify,since $\gamma_{1}\leq 1$ and $\gamma_{1}+\beta-\beta_{1}<\gamma$.
When $\gamma_{1}+\beta-\beta_{1}\geq\gamma$, we only need to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq 2+\max\{2\gamma-(\beta+1),0\}.$
We just need to check three scenarios:
(1) $2\gamma_{1}-(\beta_{1}+1)\leq 2$, since $\gamma_{1}\leq 1$.
(2) $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 2$, since $\gamma_{1}+\beta-\beta_{1}\geq\gamma\geq 0$.
(3) $2\gamma_{1}-(\beta_{1}+1)+2\gamma-(\gamma_{1}+\beta-\beta_{1})=2\gamma+\gamma_{1}-(\beta+1)\leq 2+2\gamma-(\beta+1)$, since $\gamma_{1}\leq 1$.
Case 4.2: $\gamma>\beta,\gamma_{1}\leq\beta_{1}$
In this case,
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m+1-l)(k-l+1)+\max\{2\gamma_{1}-\beta_{1},0\}.$
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k.$
$\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}.$
It suffices to prove
$\max\{2\gamma_{1}-\beta_{1},0\}\leq 1+\max\{2\gamma-(\beta+1),0\}.$
We just use $2\gamma_{1}-\beta_{1}\leq 1$, since $\gamma_{1}\leq\beta_{1}$.
Case 4.3: $\gamma\leq\beta,\gamma_{1}>\beta_{1}$
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq\begin{cases}k(n-1-k)-(m-l)(k-l+1)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}&m>l\\ k(n-1-k)&m=l.\end{cases}$
Since $\gamma_{1}+\beta-\beta_{1}>\gamma$,
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}.$
Also, we can assume $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$; otherwise, it $\infty$.
Also, we have
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
When $m=l$, we only need to prove
$\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq\max\{2\gamma-\beta,0\}.$
It is easy to verify $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 2\gamma-\beta$.
When $m>l$, it suffices to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq 1+\max\{2\gamma-\beta,0\}.$
(1) $2\gamma_{1}-(\beta_{1}+1)\leq 1+2\gamma-\beta$, since $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$ and $\gamma_{1}\leq 1$.
(2) $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 1+2\gamma-\beta$, since $\gamma_{1}>\beta_{1}$.
(3) $2\gamma_{1}-(\beta_{1}+1)+2\gamma-(\gamma_{1}+\beta-\beta_{1})=2\gamma+\gamma_{1}-(\beta+1)\leq 2\gamma-\beta$, since $\gamma_{1}\leq 1$.
Case 4.4: $\gamma\leq\beta,\gamma_{1}\leq\beta_{1}$
$\mathbf{M}(a_{1},s_{1};n-1,k)=k(n-1-k)-(m+1-l)(k-l+1)+\max\{2\gamma_{1}-\beta_{1},0\}.$
$\mathbf{M}(l+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq\begin{cases}k&\gamma>\gamma_{1}+\beta-\beta_{1}\\ k-(k-l)+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}&\gamma\leq\gamma_{1}+\beta-\beta_{1}.\end{cases}$
We can also assume $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$; otherwise, it $\infty$.
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
When $\gamma>\gamma_{1}+\beta-\beta_{1}$, it suffices to show
$\max\{2\gamma_{1}-\beta_{1},0\}\leq 1+\max\{2\gamma-\beta,0\}.$
This is true by noting $\gamma_{1}\leq 1$ and $\gamma>\gamma_{1}+\beta-\beta_{1}$.
When $\gamma\leq\gamma_{1}+\beta-\beta_{1}$, it suffices to show
$\max\{2\gamma_{1}-\beta_{1},0\}+\max\{2\gamma-(\gamma_{1}+\beta-\beta_{1}),0\}\leq 2+\max\{2\gamma-\beta,0\}.$
(1) $2\gamma_{1}-\beta_{1}\leq 2+2\gamma-\beta$, since $\gamma_{1}+\beta-\beta_{1}\leq\gamma+1$, $\gamma_{1}\leq 1$.
(2) $2\gamma-(\gamma_{1}+\beta-\beta_{1})\leq 2+2\gamma-\beta$, since $\beta_{1}\leq 1$.
(3) $2\gamma_{1}-\beta_{1}+2\gamma-(\gamma_{1}+\beta-\beta_{1})=\gamma_{1}-\beta+2\gamma\leq 2+2\gamma-\beta$, since $\gamma_{1}\leq 1$.
Case 5: $m_{1}=m-1,l_{1}=l$
Case 5.1: $\gamma>\beta,\gamma_{1}>\beta_{1}$
In this case, we have
$\mathbf{M}(m-1+\beta_{1},l+\gamma_{1};n-1,k)\leq k(n-1-k)-(m-1-l)(k-l)+\max\{2\gamma_{1}-(\beta_{1}+1),0\}.$
$\mathbf{M}(l+1+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k-(k-l)+\max\{2\gamma-(1+\gamma_{1}+\beta-\beta_{1}),0\}.$
We assume
$\gamma>\gamma_{1}+\beta-\beta_{1}.$ (81)
Otherwise, this is -$\infty$.
We also have $\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}$. We just need to prove
$\max\{2\gamma_{1}-(\beta_{1}+1),0\}+\max\{2\gamma-(1+\gamma_{1}+\beta-\beta_{1}),0\}\leq\max\{2\gamma-(\beta+1),0\}.$
Check three scenarios:
(1) $2\gamma_{1}-(\beta_{1}+1)\leq 2\gamma-(\beta+1)$, since $\gamma>\gamma_{1}+\beta-\beta_{1}$ and $\gamma_{1}\leq\gamma$ (since $s_{1}\leq s$).
(2) $2\gamma-(1+\gamma_{1}+\beta-\beta_{1})\leq 2\gamma-(\beta+1)$, since $\gamma_{1}>\beta_{1}$.
(3) $2\gamma_{1}-(\beta_{1}+1)+2\gamma-(1+\gamma_{1}+\beta-\beta_{1})=2\gamma+\gamma_{1}-(\beta+2)\leq 2\gamma-(\beta+1)$, since $\gamma_{1}\leq 1$.
Case 5.2: $\gamma>\beta,\gamma_{1}\leq\beta_{1}$
In this case, by discussing $m=l$ or $m>l$, we have
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m-l)(k-l)+\max\{2\gamma_{1}-\beta_{1},0\}.$
$\mathbf{M}(l+1+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k-(k-l)+\max\{2\gamma-(1+\gamma_{1}+\beta-\beta_{1}),0\}.$
$\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}.$
It suffices to prove
$\max\{2\gamma_{1}-\beta_{1},0\}+\max\{2\gamma-(1+\gamma_{1}+\beta-\beta_{1}),0\}\leq 1+\max\{2\gamma-(\beta+1),0\}.$
(1) $2\gamma_{1}-\beta_{1}\leq 1$, since $\gamma_{1}\leq\beta_{1}$.
(2) $2\gamma-(1+\gamma_{1}+\beta-\beta_{1})\leq 2\gamma-\beta$, since $\beta_{1}\leq 1$.
(3) $2\gamma+\gamma_{1}-(1+\beta)\leq 2\gamma-\beta$, since $\gamma_{1}\leq 1$.
Case 5.3: $\gamma\leq\beta,\gamma_{1}>\beta_{1}$
$\mathbf{M}(l+1+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)=-\infty.$
There is nothing to do.
Case 5.4: $\gamma\leq\beta,\gamma_{1}\leq\beta_{1}$
$\mathbf{M}(a_{1},s_{1};n-1,k)\leq k(n-1-k)-(m-l)(k-l)+\max\{2\gamma_{1}-\beta_{1},0\}.$
$\mathbf{M}(l+1+\gamma_{1}+\beta-\beta_{1},l+\gamma;k+1,k)\leq k-(k-l)+\max\{2\gamma-(1+\gamma_{1}+\beta-\beta_{1}),0\}.$
We assume
$\gamma>\gamma_{1}+\beta-\beta_{1}.$ (82)
Otherwise, this is $\infty.$
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
It suffices to show
$\max\{2\gamma_{1}-\beta_{1},0\}+\max\{2\gamma-(1+\gamma_{1}+\beta-\beta_{1}),0\}\leq\max\{2\gamma-\beta,0\}.$
(1) $2\gamma_{1}-\beta_{1}\leq 2\gamma-\beta$, since $\gamma>\gamma_{1}+\beta-\beta_{1}$ and $\gamma_{1}\leq\gamma$.
(2) $2\gamma-(1+\gamma_{1}+\beta-\beta_{1})\leq 2\gamma-\beta$, since $\beta_{1}\leq 1$.
(3) $2\gamma+\gamma_{1}-\beta-1\leq 2\gamma-\beta$, since $\gamma_{1}\leq 1$. ∎
Proof of Proposition 8.1.
We just combine Lemma 8.2 and Lemma 8.3. ∎
Proof of Theorem 1.11.
Finally, we see that Theorem 1.11 can be obtained from Proposition 7.1 and Proposition 8.1 by induction on $(n,k)$. ∎
Acknowledgments
The author is grateful to the anonymous reviewers for finding typos in the paper.